### v2

# Sebastian Bervoets  & Mathieu Faure  & Ludovic Renou AMSE, sebastian.bervoets(at)univ-amu.frAMSE, mathieu.faure(at)univ-amu.frQMUL and CEPR, lrenou.econ(at)gmail.com

Abstract
Deviations from Bayesian updating are traditionally categorized as biases, errors, or fallacies, thus implying their inherent “sub-optimality.” We offer a more nuanced view. We demonstrate that, in learning problems with misspecified models, non-Bayesian updating can outperform Bayesian updating.
Keywords: learning, Bayesian, consistency.
JEL Classification: C73, D82.
1 Introduction
Psychologists, economists, and other social scientists have extensively documented systematic departures from Bayesian updating. These departures include underreaction and overreaction to new information, confirmation bias, base-rate neglect, and numerous other phenomena. [^1] For comprehensive reviews, see Benjamin (2019) and Ortoleva (2022). While these departures are traditionally characterized as biases, errors, or fallacies – implying their inherent “sub-optimality” – we offer a more nuanced view. Specifically, we demonstrate that in classical learning problems, some non-Bayesian updating can outperform its Bayesian counterpart.
The benefits of non-Bayesian updating have precedent in the literature, particularly in strategic environments where such departures can induce equilibrium behavior favorable to non-Bayesian agents. In sender-receiver games, for instance, non-Bayesian updating by receivers can elicit more information from the senders (De Clippel and Zhang, 2022; Lee et al., 2023). However, these results rely fundamentally on strategic interactions. In single-agent learning environments, where strategic responses are absent, we identify a distinct mechanism through which non-Bayesian updating may confer a benefit, which we now explain.
We formalize our analysis within the canonical learning framework of Berk (1966). Consider an agent attempting to learn a true data generating process. The agent begins with a set of candidate processes, none of which may exactly match the true process. Then, in each of an infinite number of periods, the agent sequentially observes a new piece of data and updates their beliefs about which process best describes the observations. We model the agent as a conservative Bayesian, following the updating rule:
(belief at $(\text{belief at }n+1)=(1-\gamma)\times(\text{belief at }n)+\gamma\times(\text{Bayesian update at }n+1),$ (ER)
where $\gamma$ may depend on the number of observations, current belief, and current observation. This rule, first introduced by Phillips and Edwards (1966) and Edwards (1968), generalizes Bayesian updating and captures several documented cognitive biases, including under-reaction and over-reaction to new information, confirmation bias, and base-rate neglect.
In the first part of the analysis, we assume that the more observations the agent has, the less the agent reacts to new information. A body of evidence supports this assumption; in his meta-analysis, Benjamin (2019) reports it as Stylized Fact 2. Our main result – Theorem 1 – states that the agent’s predictive process – the expectation of candidate processes under the agent’s beliefs – converges almost surely to the mixture process the closest to the true data generating process (in the sense of the Kullback-Leibler divergence). In other words, the theorem states that it is as if the agent considers all possible mixtures of the candidate processes (thus, “convexifying” his model) and learns as a classical Bayesian. In particular, if the true data generating process is a mixture of the candidate processes, the agent’s predictive process converges to it. This contrasts with the Bayesian agent, whose predictive process converges to the process the closest to the true data generating process – see Berk (1966). (We stress that when the model is correctly specified, the updating rule (ER) converges to the true data-generating process, as Bayesian updating does.)
The intuition is as follows. Bayesian agents extract all the information there is from their models. [^2] Zellner (1988) proves that Bayesian updating corresponds to extracting all the information from the prior and the statistical model. However, when their models are wrong, this typically causes them to make wrong inferences. A more conservative approach, as embodied by the updating rule (ER), limits the severity of this problem. Moreover, the more observations an agent has, the more conservative they must be to counteract the increasing confidence of their Bayesian counterpart. Ultimately, this can lead to better learning outcomes.
Mathematically, we rewrite the updating rule (ER) as the Robbins-Monro algorithm:
[…]
The first term in brackets is deterministic, while the second term is random (a martingale difference). When the agent reacts sufficiently slowly to new information, the random term becomes negligible as the number of observations grows arbitrarily large. The deterministic part then drives the belief dynamics. We show that the cross entropy of the predictive process with respect to the true data generating process acts as a potential for the deterministic dynamics. [^3] The KL divergence of $p$ from $q$ is $\mathbb{E}_{p}[\log(p/q)]$. Minimizing the KL divergence with respect to $q$ is equivalent to maximizing $\mathbb{E}_{p}[\log q]$. We call the latter the cross-entropy of $q$ with respect to $p$. We then show that all critical points of the deterministic dynamics correspond to either the unconstrained maximizers of the cross-entropy or constrained ones. Finally, we prove that all local maximizers repel the dynamics, hence the dynamics converges to the unconstrained maximizers. This last step is the most delicate. The constrained maximizers are neither isolated points nor interior, which precludes us from appealing to classical results in stochastic approximation theory such as Pemantle (1990), Brandiere (1998), or Brandiere and Duflo (1996). We develop new arguments to solve that issue.
In the second part of the analysis, we consider alternative assumptions on the updating weights and show that our main insight, that is, that some non-Bayesian updating can outperform Bayesian updating, continues to hold. For instance, if the updating weights are constant, then the belief dynamics does not converge, but its occupation measure does. As the weight $\gamma$ becomes arbitrarily small, the occupation measure concentrates on the mixture process the closest to the true data generating process – Theorem 3.
We illustrate our main result with a simple example, inspired by the work of Spiegler (2016) on causal models. Suppose that each observation is a vector of variables $(x^{1},x^{2},x^{3})$, representing diet, sleep, and health, respectively. The true causal model is such that diet and sleep function are interdependent determinants of health, with diet also influencing sleep patterns. However, the agent posits another causal model and assume that diet and sleep are independent causes of health. Figure 1 represents the causal models as two directed acyclic graphs. Thus, the agent considers distributions of the form $p(x^{3}|x^{1},x^{2})p(x^{2})p(x^{1})$, which all differ from the true distribution since $x^{1}$ and $x^{2}$ are correlated in the true model. Since any correlated distribution is the convex combination of independent distributions, our main result implies that the non-Bayesian agent successfully converges to the correct causal model, while the Bayesian agent fails to do so. This is the paper’s main insight.
[…]
Figure 1: The causal models
We also contrast our results with the generalized Bayes rule
$(\text{belief at }n+1)\propto(\text{belief at }n)\times(\text{Bayesian update at }n+1)^{\gamma},$ (GR)
a special case of Grether rule (Grether, 1980) and show that its behavior is fundamentally different. While also underreacting to new information, the updating rule (GR) has the same asymptotic properties than Bayesian updating. Underreaction alone does not explain our results.
We close the introduction with a brief discussion of the related literature (Section 5 contains a detailed discussion). The two most closely related papers are Berk (1966) and Epstein et al. (2010). Berk (1966) studies the long-run properties of Bayesian updating when the model is misspecified and shows that the beliefs converge to the process, which minimizes the KL divergence from the true process.444 See also Freedman (1963), Berk (1970), Diaconis and Freedman (1986), Diaconis and Freedman (1999), Shalizi (2009). Epstein et al. (2010) study the long-run properties of the updating rule (ER) when the model is correctly specified. They show that if the agent under-reacts to new information, the agent eventually learns the true process. We differ from Berk (1966) by assuming non-Bayesian updating and from Epstein et al. (2010) by assuming a misspecified model. Some recent papers, e.g., Bohren and Hauser (2021, 2024), Heidhues et al. (2021) and Frick et al. (2023), provide a link between Bayesian updating in misspecified problems and non-Bayesian updating. In particular, we can construct another set of candidate processes such that Bayesian updating in that modified model induces the same beliefs than the updating rule (ER) in the original model. To the best of our understanding, none of these recent results are applicable to our problem.
2 The Problem
At each period $n\in\mathbb{N}$, an agent is observing $x_{n}\in X$, independently and identically drawn from an unknown data-generating process $p^{*}\in\Delta(X)$. The agent postulates a family $\mathcal{P}:=\{p_{\theta}\in\Delta(X):\theta\in\Theta\}$ of finitely many possible data-generating processes, which may not include the true data-generating process $p^{*}$ – in that case, the model is misspecified. Throughout, we assume that the sets $X$ and $\Theta$ are finite and that the support of $p^{*}$ and each $p_{\theta}$ is $X$ [^5] None of our results would change if we were to assume that the support of $p^{*}$ is included in the support of $p_{\theta}$ for all $\theta\in\Theta$. The only modification would be to substitute $X$ with the support of $p^{*}$ in the analysis. The latter assumption guarantees that the agent is never surprised, that is, the agent cannot observe a realization $x$, believed impossible under all data-generating processes.
We think of $x$ as financial returns in investment problems, losses in insurance problems, or demands in pricing problems. For instance, in a canonical portfolio problem, $x$ models the realized asset returns and $\mathcal{P}$ their postulated distributions, e.g., multivariate Gaussian distributions with parameters $\theta$, i.e., means and variance-covariance matrices. As another instance, in canonical insurance problems, $x$ models the losses or, equivalently, the final wealth. More generally, we view $x$ as payoff-relevant consequences, $p^{*}$ as their true distribution, and $\mathcal{P}$ as the collection of “parameterized” distributions the agent postulates. An alternative is to view $x$ as a (payoff-irrelevant) signal about the (payoff-relevant) state $\theta$, $p^{*}$ as the signal distribution when the true state is $\theta^{*}\in\Theta$, and $(p_{\theta})_{\theta}$ as the perceived conditional distributions. Both views are common in Economics and Statistics. While our formal results are independent of either perspective, their interpretations differ. Under the first view, successful learning entails the convergence of the predictive process to the true distribution $p^{*}$, whereas under the alternative view, learning success is characterized by the identification of the true state $\theta^{*}$.
The agent’s prior is $q_{0}\in\Delta(\Theta)$. Without loss of generality, we assume that the support of $q_{0}$ is $\Theta$, i.e., $q_{0}\in\mathbf{S}^{*}:=\mathrm{int}\Delta(\Theta)$, the interior of $\Delta(\Theta)$. The agent updates his belief from $q_{n}$ to $q_{n+1}$ upon observing the signal $x_{n+1}$ as follows:
$q_{n+1}=(1-\gamma_{n+1})q_{n}+\gamma_{n+1}B(q_{n},x_{n+1}),$ (ER)
where $B(q_{n},x_{n+1})$ is the Bayesian posterior of $q_{n}$ given the new observation $x_{n+1}$, and $\gamma_{n+1}$ the (possibly random and positive) weight on the Bayesian posterior. The updating rule (ER) was first discussed in Phillips and Edwards (1966) and Edwards (1968), generalizes the Bayesian rule ($\gamma_{n+1}=1$) and can accommodate a number of well-documented biases such as underreaction ($\gamma_{n+1}<1$) and overreaction ($\gamma_{n+1}>1$) to new information, confirmatory biases, and a few others. We refer the reader to Epstein (2006) and Kovach (2021) for more extensive discussions and axiomatic foundations. Benjamin (2019) documents two additional empirical findings, which the rule can also accommodate. First, the more observations an agent has, the more the agent under-reacts. This finding – Stylized Fact 2 in Benjamin (2019) – is consistent with decreasing updating weights $(\gamma_{n})_{n}$. Second, individuals’ beliefs after sequentially observing $x_{1}$ to $x_{n}$ differ from the beliefs they would have formed, had they received the same information simultaneously as a single observation $(x_{1},\dots,x_{n})$. This finding – Stylized Fact 8 in Benjamin (2019) – is also consistent with the updating rule (ER), even if we assume constant updating weights. [^6] Cripps (2018) characterizes the updating rules, which produce the same update after observing either $x_{1}$ to $x_{n}$ sequentially, or $(x_{1},\dots,x_{n})$ simulatenously
We analyze the long-run properties of the agent’s predictive process, i.e., $(\mathbb{E}_{q_{n}}[p_{\bm{\theta}}])_{n\in\mathbb{N}}$, under updating rule (ER), with a particular focus on prediction accuracy. While a frequentist approach would guarantee correct learning in this setting (a direct application of the law of large numbers), its applicability is limited in many other settings. For instance, in all settings where the agent observes a single realization, the frequentist approach has no bite. As a generalization of Bayesian updating, the updating rule (ER) does not suffer from that issue. [^7] E.g., Lee et al. (2023) revisit the canonical cheap-talk model of Crawford and Sobel (1982) and show that the receiver can be better off using rule (ER). We focus on the updating rule (ER) for its applicability to a wide range of settings and ability to explain empirical regularities on belief updating.
We conclude this section with a brief review of Berk (1966) and Epstein et al. (2010). Berk (1966) studies the convergence of the Bayesian posteriors $(q_{n})_{n}$ and proves the almost-sure convergence to the maximizers of the (negative of the) cross entropy, that is, $\arg\max_{\theta\in\Theta}\sum_{x}p^{*}(x)\log p_{\theta}(x)$. Epstein et al. (2010) study whether the updating rule (ER) induces the agent to eventually learn the true data-generating process, when the model is correctly specified. They show that if the agent is always underreacting to new information (i.e., $\gamma_{n+1}\leq 1$), then the agent eventually learns the true data-generating process, when the weight $\gamma_{n+1}$ is measurable with respect to the information up to period $n$. With the help of an example, they show that this result is not true when the weight $\gamma_{n+1}$ is measurable with respect to the information up to period $n+1$.
3 Underreaction and Learning
This section presents our main result. We assume that $(\gamma_{n})_{n}$ is a deterministic sequence of positive real numbers such that
$\sum_{n}\gamma_{n}=+\infty,\text{ and }\sum_{n}\gamma^{2}_{n}<+\infty.$ (1)
E.g., the series $(1/n)_{n}$ satisfies the condition. Intuitively, the more observations the agent has, the more the agent under-reacts to new information ($\sum_{n}\gamma^{2}_{n}<+\infty$). Yet, the underreaction is not so severe that the agent stops reacting ($\sum_{n}\gamma_{n}=+\infty$). Stylized Fact 2 in Benjamin (2019) motivates the assumption. We discuss several alternatives in the next section and show that the corresponding characterizations are qualitatively the same as the one we now present.
We observe that the updating rule (ER) can be rewritten as:
$q_{n+1}-q_{n}=\gamma_{n+1}\left(H(q_{n})+U_{n+1}\right),$
where $H:\Delta(\Theta)\rightarrow\mathbb{R}_{0}^{|\Theta|}$ is the multi-valued function, whose $\theta$-th coordinate is
$H_{\theta}(q):=q_{\theta}\Biggl{(}\underbrace{\sum_{x}p^{*}(x)\frac{p_{\theta}(x)}{\sum_{\theta^{\prime}}q_{\theta^{\prime}}p_{\theta^{\prime}}(x)}}_{:=f_{\theta}(q)}-1\Biggr{)}=\underbrace{\mathbb{E}_{p^{*}}\left[B_{\theta}(q,\mathbb{x})\right]}_{\text{Expectation of Bayesian posterior of $\theta$}}-\;q_{\theta},$ (2)
and $(U_{n})_{n}$ a bounded martingale difference, defined by $U_{n+1}:=B(q_{n},x_{n+1})-\mathbb{E}_{p^{*}}[B(q_{n},\mathbb{x}_{n+1})]$. [^8] The very same arguments apply if the agent updates his belief from $q_{n}$ to $q_{n+1}:=(1-\gamma_{n+1})q_{n}+\gamma_{n+1}b(q_{n},x_{n+1})$, where $b(q_{n},x_{n+1})$ is a noisy estimate of $B(q_{n},x_{n+1})$ such that $\mathbb{E}[b(q,x)|B(q,x)]=B(q,x)$ for all $(q,x)$. This recursive formulation is a Robbins-Monro algorithm (Robbins and Monro, 1951), a well-known algorithm in the literature on stochastic approximation theory. A popular method for analyzing such algorithms is the ordinary differential equation (ODE) method, which approximates the sequence $(q_{n})_{n}$ with the solution to the associated (deterministic) ODE:
$\dot{q}(t)=H(q(t)).$
When applicable, this method reduces the analysis of the Robbins-Monro algorithm (ER) to the study of the associated ODE. We prove two central results about the ODE. (All proofs are in the Appendix.) The first result – Lemma 1 – states that the function $q\mapsto V(q)$ is a Lyapunov function for the ODE, where $V(q)$ is the (negative of the) cross-entropy of the mixture $\sum_{\theta}q_{\theta}p_{\theta}$ with respect to $p^{*}$, that is,
$V(q):=\sum_{x}p^{*}(x)\log\left(\sum_{\theta}q_{\theta}p_{\theta}(x)\right).$ (3)
The second result – Lemma 2 – states that the set $E:=\left\{q\in\Delta(\Theta):\;\,H(q)=0\right\}$ of critical points of $H$ can be decomposed into a finite union of non-empty, disjoint, compact and convex components $C_{k}$, $k=1,\dots,K$, with the cross-entropy $V$ constant on each of the components. (Note that the set of degenerated beliefs $\{\delta_{\theta}:\theta\in\Theta\}$ belong to $E$, so that $E$ is non-empty.) The decomposition is unique. For further reference, we denote $C_{k^{*}}$ the unique component with the maximal value of $V$.
Definition 1
The component $C_{k^{*}}$ is $\{q\in E:V(q)\geq V(q^{\prime}),\forall q^{\prime}\in E\}$, with $E$ the set of critical points of $H$.
[…]
Example 1
Assume that $\Theta=\{\theta_{1},\theta_{2},\theta_{3}\}$, $X=\{x_{1},x_{2},x_{3}\}$, and
$p_{\theta_{1}}=\frac{1}{5}(1,2,2),\;p_{\theta_{2}}=\frac{1}{5}(2,1,2),\;p_{\theta_{3}}=\frac{1}{5}(2,2,1),\,p^{*}=\frac{1}{3}(1,1,1).$
The set $E$ is:
$\left\{\delta_{\theta_{1}}\right\}\cup\left\{\delta_{\theta_{2}}\right\}\cup\left\{\delta_{\theta_{3}}\right\}\cup\left\{\frac{1}{2}(1,1,0)\right\}\cup\left\{\frac{1}{2}(1,0,1)\right\}\cup\left\{\frac{1}{2}(0,1,1)\right\}\cup\left\{\frac{1}{3}(1,1,1)\right\}.$
We have $V(\delta_{\theta_{1}})=V(\delta_{\theta_{2}})=V(\delta_{\theta_{3}})=\frac{1}{3}\log\frac{1}{5}+\frac{2}{3}\log\frac{2}{5}$,
$V\left(\frac{1}{2}(1,1,0)\right)=V\left(\frac{1}{2}(1,0,1)\right)=V\left(\frac{1}{2}(0,1,1)\right)=\frac{2}{3}\log\frac{3}{10}+\frac{1}{3}\log\frac{4}{10},$
and $V\left(\frac{1}{3}(1,1,1)\right)=\log\frac{1}{3}$. Here, $C_{k^{*}}=\left\{\frac{1}{3}(1,1,1)\right\}$, the interior point.
$\theta_{3}$$\theta_{1}$$\theta_{2}$
Figure 2: The components in Example 1
To sum up, the cross-entropy $V$ is a Lyapunov function for the flow of the dynamics and maximized in a unique component $C_{k^{*}}$ of critical points. This suggests that the component $C_{k^{*}}$ is globally stable. While our main result – Theorem 1 – states that this is indeed the case, the result is far from immediate. As we shall see, the main technical challenge is that the components $C_{k}$ are not always singletons (isolated points) and, moreover, often exist as continuums on the boundary.
Theorem 1
For all $q\in\mathrm{int}\Delta(\Theta)$, the (random) limit set $\mathcal{L}((q_{n})_{n})$ of the process $(q_{n})_{n}$ is included in $C_{k^{*}}$ $p^{*}$-almost surely, that is, any limit process is a maximizer of the cross-entropy $V$.
Theorem 1 thus states that every $\omega$-limit $q$ induces a unique distribution $\sum_{\theta}q_{\theta}p_{\theta}\in{\rm co\,}\mathcal{P}$, which attains the maximum:
$\max_{p\in{\rm co\,}\mathcal{P}}\sum_{x}p^{*}(x)\log p(x).$
It follows immediately that:
Corollary 1
If $p^{*}\in{\rm co\,}\mathcal{P}$, then the agent’s predictive process converges to the true data-generating process $p^{*}$.
Thus, if $p^{*}\in{\rm co\,}\mathcal{P}$, the learning outcome is the same as the one of a Bayesian learner with the richer set of (non-parametric) models ${\rm co\,}\mathcal{P}$, assuming that the agent’s prior assigns a positive probability to an open neighborhood of the cross-entropy maximizer – see Berk (1966). This observation is reminiscent of Bohren and Hauser (2024), who demonstrate an equivalence between non-Bayesian updating in a correctly specified model and Bayesian updating in a misspecified model. In particular, if the agent’s model is $\{\pi_{\theta}^{q_{n},\gamma_{n}}:\theta\in\Theta\}$, with
$\pi_{\theta}^{q_{n},\gamma_{n}}(x):=(1-\gamma_{n})\mathbb{E}_{q_{n}}[p_{\bm{\theta}}(x)]+\gamma_{n}p_{\theta}(x)$
for all $x$, then the Bayesian posteriors of the modified model $\{\pi_{\theta}^{q_{n},\gamma_{n}}:\theta\in\Theta\}$ coincide with the non-Bayesian posteriors (ER) of the original model $\{p_{\theta}:\theta\in\Theta\}$, when the agent’s prior belief is $q_{n}$ at period $n$. We stress that the “as-if” model depends on the current belief $q_{n}$ and weight $\gamma_{n}$. The dependence on current beliefs is a common feature of social learning models, see e.g., Esponda and Pouzo (2016), Bohren and Hauser (2021), and Frick et al. (2023). However, as far as we know, none of the results developed in these (and other related) papers are directly applicable to our problem. Indeed, the main focus of this literature is on the stability of degenerated beliefs and none of the few instances, where mixed beliefs are considered, are directly applicable in our setting. See Section 5 for a more detailed discussion.
Thus, conservative updating protects the agent against misspecification and leads to better predictions. Predicting better, even perfectly, does not imply that the agent learns (identifies) the true data-generating process. In fact, the agent cannot learn the true data-generating process when his model is misspecified. Yet, if we interpret $\Theta$ as states and $X$ as signals, an implication of Theorem 1 is that the agent’s belief about the states may also come closer to the true state than Bayesian updating would. For an illustration, suppose that there are two states $\theta^{*}$ and $\theta$ and two signals $\ell$ and $h$. The true conditional distributions (experiment) are $p^{*}_{\theta^{*}}(h)=1-p^{*}_{\theta}(h)=2/3$, while the agent’s model is $p_{\theta^{*}}(h)=1-p_{\theta}(h)=1/10$. Assume that the true state is $\theta^{*}$. Since the agent wrongly interprets the signal $h$ as evidence of $\theta$ and the signal $h$ is the most likely when the state is $\theta^{*}$, the Bayesian posterior converges almost surely to probability one on $\theta$, the wrong state. In contrast, with the updating rule (ER), the agent’s belief converges to probability $7/24$ on $\theta^{*}$, a closer distribution to the truth. (Its cross-entropy with respect to $\delta_{\{\theta^{*}\}}$ is $\log(7/24)$, while the cross-entropy of the limiting Bayesian posterior is $\infty$- ∞. [^9] We compare distributions in $\Delta(\Theta)$. Since the parameter identifies distributions in $\Delta(X)$, we are equivalently comparing distributions in $\Delta(\Delta(X))$.) It is not always true that the non-Bayesian agent learns better, though. E.g., if we change $1/10$ for $9/10$, the agent learns the true state under Bayesian updating, while the belief converges to $17/24$ on $\theta^{*}$ under updating (ER). In general, all we can say is that the agent’s belief comes closer, and sometimes strictly so, to the true state under updating (ER) (in the sense of the Kullback-Leibler divergence) if, and only if, the agent does not learn the true state under Bayesian updating.
The updating rule (ER) may also serve as a “misspecification test.” Indeed, if the process $(q_{n})_{n}$ does not converge to a point mass on one of the parameter $\theta$, then the model $\mathcal{P}$ must be misspecified. [^10] More precisely, we need to add the mild requirement that the model is identified, i.e., $p_{\theta}\neq p_{\theta^{\prime}}$ for all $(\theta,\theta^{\prime})$. When $p^{*}=p_{\theta}=p_{\theta^{\prime}}$ with $\theta\neq\theta^{\prime}$, it is possible for the belief process to converge to an interior point, even though the model is correctly specified. Thus, a possible test is to reject the hypothesis that the model is correctly specified if $\min_{\theta\in\Theta}||q_{n}-\delta_{\{\theta\}}||_{TV}>\varepsilon$, where $\varepsilon$ and $n$ are carefully calibrated to trade-off type I and II errors. We leave this open question for future research.
The proof of Theorem 1 uses results from stochastic approximation theory with, however, one major difficulty. The major difficulty is that the convex components of $E$ might neither be isolated points nor interior points. In fact, each component $C_{k}\neq C_{k^{*}}$ lies on the boundary of $\Delta(\Theta)$. Example 2 illustrates that we may have a continuum of critical points on the boundary.
Example 2
Suppose that $\Theta=\{\theta_{1},\theta_{2},\theta_{3},\theta_{4}\},X=\{x_{1},x_{2},x_{3}\}$, and
$p_{\theta_{1}}=\frac{1}{4}(2,1,1),\;p_{\theta_{2}}=\frac{1}{4}(1,2,1),\;p_{\theta_{3}}=\frac{1}{4}(1,1,2),\;p_{\theta_{4}}=\frac{1}{8}(2,3,3),p^{*}=\frac{1}{8}(6,1,1).$
The set $E$ is
$\{\delta_{\theta_{1}}\}\cup\{\delta_{\theta_{2}}\}\cup\{\delta_{\theta_{3}}\}\cup\{(0,\lambda/2,\lambda/2,1-\lambda):\;\lambda\in[0,1]\},$
with $C_{k^{*}}=\{\delta_{\theta_{1}}\}$. The component $\{(0,\lambda/2,\lambda/2,1-\lambda):\;\lambda\in[0,1]\}$ is a continuum and lies on the boundary of $\Delta(\Theta)$ – it is the line in Figure 3.
$\theta_{1}$$\theta_{3}$$\theta_{4}$$\theta_{2}$
Figure 3: The components in Example 2
As a consequence, none of the classical results in stochastic approximation theory, e.g., Pemantle (1990), Brandiere (1998), Brandiere and Duflo (1996), are immediately applicable. For instance, Pemantle (1990, p. 700-701) states that “if there are uncountably many unstable points, then $(q_{n})$ can still converge to one of them even though each point has probability zero of being hit. In this case, all one might conclude is that the law of the limit has no point masses at unstable critical points.” Similarly, since the components $C_{k}$, $k\neq k^{*}$, are on the boundary of $\Delta(\Theta)$, we cannot rely on the Jacobian of $H$ in the neighborhood of each $q\in C_{k}$ to determine their stability.
We develop novel arguments to tackle that issue. The arguments are quite technical, but the main logic is simple: We show that each component $C_{k}\neq C_{k^{*}}$ admits an open neighborhood and a uniform repelling direction, so that the dynamics cannot hop indefinitely often from one element of the open neighborhood of $C_{k}$ to another.
We now present the arguments for the special case where $C_{k}\neq C_{k^{*}}$ consists of the single element ${\hat{q}}$. (Technical details in this paragraph may be omitted without loss of continuity.) Choose $q^{*}\in C_{k^{*}}$ with maximal support, that is, for all $q\in C_{k^{*}}$, $[q_{\theta}>0\Rightarrow q^{*}_{\theta}>0]$. The convexity of each component guarantees that such $q^{*}$ exists. We partition $\Theta$ into three subsets: $\Theta_{1}:=\{\theta\in\Theta:\;q^{*}_{\theta}=\hat{q}_{\theta}=0\}$, $\Theta_{2}:=\{\theta\in\Theta:\;\hat{q}_{\theta}>0\}$, and $\Theta_{3}:=\left\{\theta\in\Theta:\;\hat{q}_{\theta}=0,q^{*}_{\theta}>0\right\}$. From the construction of the components $C_{k}$, $\Theta_{3}$ is non-empty. (If $\Theta_{3}$ was empty, then $C_{k}=C_{k^{*}}$.) From the strict concavity of $V$, $\frac{\partial}{\partial\tau}V\left(\tau q^{*}+(1-\tau)\hat{q}\right)>0$. Therefore,
$0<\sum_{\theta\in\Theta}(q^{*}_{\theta}-\hat{q}_{\theta})f_{\theta}(\hat{q})=\sum_{\theta\in\Theta}q^{*}_{\theta}f_{\theta}(\hat{q})-1=\sum_{\theta\in\Theta_{2}\cup\Theta_{3}}q^{*}_{\theta}f_{\theta}(\hat{q})-1=\sum_{\theta\in\Theta_{2}}q^{*}_{\theta}+\sum_{\theta\in\Theta_{3}}q^{*}_{\theta}f_{\theta}(\hat{q})-1,$
where the first and third equalities follows from $\hat{q}_{\theta}(f_{\theta}(\hat{q})-1)=0$ for all $\theta$, since $\hat{q}$ is a critical point of $H$. Since $\sum_{\theta\in\Theta_{2}\cup\Theta_{3}}q_{\theta}^{*}=1$, there exists $\hat{\theta}\in\Theta_{3}$ such that $f_{\hat{\theta}}(\hat{q})>1$, that is, $\dot{q}_{\hat{\theta}}>0$ in the neighborhood of $\hat{q}$ – the dynamics cannot converge to $\hat{q}$. The logic extends to arbitrary convex components.
The observation that the components are not necessarily isolated points is a direct consequence of the fact that distributions in ${\rm co\,}\mathcal{P}$ do not always have a unique decomposition in terms of its extreme points $\mathcal{P}$. For future reference, let $\Lambda$ be the set of all possible convex combinations of the $p_{\theta}$’s equal to $p^{*}$, that is,
$\Lambda:=\left\{\lambda=(\lambda_{\theta})_{\theta\in\Theta}\in\Delta(\Theta):\;\,p^{*}=\sum_{\theta\in\Theta}\lambda_{\theta}p_{\theta}\right\}.$
We now discuss the implications of assuming different notions of independences of the family $\mathcal{P}$. We say that the family $\mathcal{P}=\{p_{\theta}:\theta\in\Theta\}$ is full if the vectors $(p_{\theta}-p_{\theta^{\prime}})_{(\theta,\theta^{\prime})\in\Theta\times\Theta}$ span the tangent space associated to $\Delta(X)$, i.e., $\mathbb{R}^{|X|}_{0}:=\{v\in\mathbb{R}^{|X|}:\,\;\sum_{x\in X}v_{x}=0\}$. In words, if the family is rich enough, then differences between its members can point in any direction allowed by the simplex structure. This property is common in statistics, as it allows for identification.
Proposition 1
If the family $\mathcal{P}$ is full, then $E\cap\mathbf{S}^{*}\subseteq\Lambda$. If, in addition, $p^{*}\notin{\rm co\,}\mathcal{P}$, then $E\subseteq\partial\left(\Delta(\Theta)\right)$.
Thus, when the family is full, interior critical points, if they exist, “identifies” the true data-generating process $p^{*}$; the agent would learn to predict correctly. We stress that the inclusion $E\cap\mathbf{S}^{*}\subseteq\Lambda$ may be strict, because $\Lambda$ may intersect the boundary of $\Delta(\Theta)$, as the following example demonstrates.
Example 3
Suppose that $\Theta=\{\theta_{1},\theta_{2},\theta_{3}\},X=\{x_{1},x_{2}\}$, and
$p_{\theta_{1}}=\frac{1}{4}(1,3);\;p_{\theta_{2}}=\frac{1}{3}(1,2);\;p_{\theta_{3}}=\frac{1}{4}(3,1),p^{*}=\frac{1}{2}(1,1).$
Then, $E=\{\delta_{\theta_{1}}\}\cup\{\delta_{\theta_{2}}\}\cup\{\delta_{\theta_{3}}\}\cup\Lambda$, where
$\Lambda=\left\{\lambda\in\Delta(\Theta):\,\lambda_{1}p_{\theta_{1}}+\lambda_{2}p_{\theta_{2}}+\lambda_{3}p_{\theta_{3}}=p^{*}\right\}=\left\{\left(\lambda_{1},\frac{1}{5}(3-6\lambda_{1}),\frac{1}{5}(2+\lambda_{1})\right):\;\lambda_{1}\in\left[0,\frac{1}{2}\right]\right\}.$
$\theta_{3}$$\theta_{1}$$\theta_{2}$
Figure 4: The four components in Example 3
We say that the family $\mathcal{P}$ satisfies convex independence if there does not exist $\hat{\theta}\in\Theta$ such that $p_{\hat{\theta}}\in{\rm co\,}\{p_{\theta}:\theta\in\Theta\setminus\{\hat{\theta}\}\}$. The next result states that all components are singletons when the family satisfies convex independence and $|\Theta|\ \leq|X|$.
Proposition 2
If $|\Theta|\leq|X|$ and the family $\mathcal{P}$ satisfies convex independence, then every $C_{k}$ is a singleton.
Finally, we say that the family $\mathcal{P}$ is tight if it is full and minimal, in the sense that any proper sub-family is not full. Note that a family is tight if and only if $|\Theta|=|X|$ and it satisfies convex independence. We have the following theorem when the family is tight.
Theorem 2
Assume that the family $\mathcal{P}$ is tight. There exists $\lambda^{*}$ such that $(q_{n})_{n}$ converges to $\lambda^{*}$ almost surely. If $\Lambda\neq\emptyset$, then $\Lambda=\{\lambda^{*}\}$.
4 Discussion
4.1 Updating Weights
Section 3 assumes that the updating weights $(\gamma_{n})_{n}$ are deterministic, square summable, but not summable. In line with some empirical evidences (Benjamin, 2019), these assumptions guarantees that the agent does not stop learning, but does so at a progressively slower rate. This section discusses alternative assumptions and their consequences.
We start with two immediate observations. First, since $||U_{n+1}||$ is bounded by 2, the updating process is sub-Gaussian. From Proposition 4.4. of Benaïm (1999), Theorem 1 remains true with the weaker condition: $\sum_{n}e^{-c/\gamma_{n}}<+\infty$ for some positive constant $c$. E.g., the series $(1/\sqrt{n})_{n\in\mathbb{N}}$ satisfies this condition (since $\sum_{n}e^{-\sqrt{n}}<+\infty$), but not the one from Section 3. Second, the updating weights can be random. E.g., Theorem 1 remains true if each random weight $\bm{\gamma}_{n}$ is measurable with respect to $\mathcal{F}_{n-1}$, $\mathbb{E}[\sum_{n}\bm{\gamma}_{n}]=+\infty$, and $\mathbb{E}[\sum_{n}\bm{\gamma}_{n}^{2}]<+\infty$.111111 We denote $(\mathcal{F}_{n})_{n}$ the filtration adapted to the sequence $(q_{n},x_{n})_{n}$. See Remark 4.3 of Benaïm (1999). We now turn our attention to two more substantial alternatives.
Constant Weights. Assume that the updating weights are constant, i.e., $\gamma_{n}=\gamma\in(0,1)$ for all $n$. With constant weights, the agent continues to underreact to information, but does so at a constant rate. The analysis in Epstein et al. (2010) covers this case and shows that the agent eventually learns the data-generating process, that is, if $p_{\theta^{*}}$ is the true data-generating process, the agent’s beliefs converge almost surely to $\delta_{\{\theta^{*}\}}$. The argument is standard. Under the true data-generating process $p_{\theta^{*}}$, the random process $(\log q_{n}(\theta^{*}))_{n}$ is a sub-martingale, bounded from above, and thus converges almost surely by the martingale convergence theorem. To see this, note that
$\mathbb{E}_{p_{\theta^{*}}}[\log q_{n+1}(\theta^{*})]-\log q_{n}(\theta^{*})$ $=\mathbb{E}_{p_{\theta^{*}}}\left[\log\left((1-\gamma)+\gamma\frac{p_{\theta^{*}}(x)}{\sum_{\theta^{\prime}}q_{n}(\theta^{\prime})p_{\theta^{\prime}}(x)}\right)\right]$
$\geq\gamma\sum_{x}p_{\theta^{*}}(x)\log\left(\frac{p_{\theta^{*}}(x)}{\sum_{\theta^{\prime}}q_{n}(\theta^{\prime})p_{\theta^{\prime}}(x)}\right)\geq 0,$
where the first inequality follows from the concavity of $\log$ and the second from the positivity of the relative entropy. It follows that $(q_{n}(\theta^{*}))_{n}$ converges almost surely to one. However, when the model is misspecified (i.e., $p^{*}\notin\mathcal{P}$), the argument does not apply. Indeed, when the model is misspecified, the term
$\gamma\sum_{x}p^{*}(x)\log\left(\frac{p_{\theta^{*}}(x)}{\sum_{\theta^{\prime}}q_{\theta^{\prime}}p_{\theta^{\prime}}(x)}\right),$
is not always positive. For some $q$, the predictive distribution $\mathbb{E}_{q}[p_{\bm{\theta}}]$ is the closest to $p^{*}$, in which case it is strictly negative.
In fact, the belief process is not even guaranteed to converge. For an example, suppose that $X=\{0,1\}$, $\Theta=\{\theta,\theta^{\prime}\}$, $p^{*}(1)=2/3$, $p_{\theta}(1)=3/4=p_{\theta^{\prime}}(0)$. We plot the belief processes in Figure 5. The orange curve represents the predictive processes induced by the updating rule (ER) with a constant weight, while the black curve represents the true data generating process. Clearly, the process does not converge, but oscillates around the true data-generating process. For completeness, we also plot the predictive processes induced by the updating rule (ER) with decreasing weights (resp., Bayesian updating) as the green (resp., purple) curve.
Figure 5: Constant weight and non-convergence
We therefore need to follow a different approach than in the previous section. We start by noting that the updating rule (ER) induces a Markov chain on $\Delta(\Theta)$, where the probability to transition from $q$ to $q^{\prime}$ is $\sum_{x\in X_{q,q^{\prime}}}p^{*}(x)$, $X_{q,q^{\prime}}:=\{x\in X:q^{\prime}=(1-\gamma)q+\gamma B(q,x)\}$. From any $q$, the Markov chain transitions to finitely many $q^{\prime}$. We then study the long-run properties of the Markov chain, which hold for almost all possible realizations of the observations. More precisely, we study the convergence properties of the sequences of occupation measures $(\Pi^{\gamma}_{n})_{n}$, where for any Borel set $A\subset\Delta(\Theta)$,
$\Pi^{\gamma}_{n}(A):=\frac{1}{n}\sum_{m=1}^{n}\delta_{A}(q_{m}).$
The notation stresses that the chain and, therefore, its occupation measure is parameterized by $\gamma$. Also, recall that $q_{m}$ is random, as it depends on the realized observations.
We first prove that the sequences of occupation measures converge to invariant distributions $\pi\in\Delta(\Delta(\Theta))$ of the Markov chains. This result is, however, of limited value as there are multiple invariant measures. For instance, for each $\theta\in\Theta$, $\delta_{\{\delta_{\theta}\}}$ is an invariant measure. We, however, prove that as $\gamma$ becomes arbitrarily small, the occupation measure converges to an invariant measure, which concentrates on $C_{k^{*}}$. (Recall that $C_{k^{*}}$ is the component of critical points, which maximize the cross-entropy.) Thus, when the reaction to new information is small enough, the process is qualitatively similar to the process with decreasing weights studied in Section 3. We now turn to a formal definition and statement.
Definition 2
The measure $\pi^{*}\in\Delta(\Delta(\Theta))$ is a limiting measure for the updating rule (ER) if there exists a sequence $(\gamma_{\ell},\pi_{\ell})_{\ell\in\mathbb{N}}$ such that:
  1. 1.
    $\gamma_{\ell}\downarrow 0$,
  2. 2.
    $\pi_{\ell}$ is a weak* limit point of $(\Pi_{n}^{\gamma_{\ell}})_{n}$, for all $\ell$,
  3. 3.
    $\lim_{\ell}\pi_{\ell}=\pi^{*}$ for the weak* topology.
In words, the measure $\pi^{*}$ is a limiting measure if we can find a sequence of updating weights $(\gamma_{\ell})_{\ell}$ and a corresponding sequence $(\pi_{\ell})_{\ell}$ of limit points of $(\Pi_{n}^{\gamma_{\ell}})_{n}$, which converges to $\pi^{*}$ for the weak* topology. We then prove the following:
Theorem 3
Let $\pi^{*}$ be a limiting measure for the updating rule (ER). The support of $\pi^{*}$ is included in $C_{k^{*}}$.
Theorem 3 states that the limiting measure concentrates on $C_{k^{*}}$ as the updating weight becomes smaller, the natural analogue of Theorem 1. To get some intuition, recall that a limit point $\pi_{\ell}$ of $(\Pi_{n}^{\gamma_{\ell}})_{n}$ is an invariant measure of the Markov chain, parameterized by $\gamma_{\ell}$. Since invariant measures capture the long-run stationary properties of the belief process, beliefs should not grow at the limit. More precisely, let
$r_{\theta}^{\gamma}(q):=\sum_{x}p^{*}(x)\underbrace{\log\left((1-\gamma)+\gamma\frac{p_{\theta}(x)}{\sum_{\theta^{\prime}}q_{\theta^{\prime}}p_{\theta^{\prime}}(x)}\right)}_{\approx\frac{q_{n+1}(\theta|x)-q_{n}(\theta)}{q_{n}(\theta)}}.$
be the expected growth rate of the belief in $\theta$, when the current belief is $q$. We show that
$\int_{\Delta(\Theta)}\sum_{\theta}\lambda_{\theta}r_{\theta}^{\gamma_{\ell}}(q)\pi_{\ell}(dq)\leq 0,$ (4)
for all positive weights $\lambda\in\mathbb{R}_{+}^{|\Theta|}$, for any limit point $\pi_{\ell}$. Now, choose $\lambda^{*}\in C_{k^{*}}$. For any $q\notin C_{k^{*}}$, we have that
$\sum_{x}p^{*}(x)\left(\frac{\sum_{\theta^{\prime}}\lambda_{\theta^{\prime}}^{*}p_{\theta^{\prime}}(x)}{\sum_{\theta^{\prime}}q_{\theta^{\prime}}p_{\theta^{\prime}}(x)}-1\right)\geq\sum_{x}p^{*}(x)\log\left(\frac{\sum_{\theta^{\prime}}\lambda_{\theta^{\prime}}^{*}p_{\theta^{\prime}}(x)}{\sum_{\theta^{\prime}}q_{\theta^{\prime}}p_{\theta^{\prime}}(x)}\right)>0,$
where the first inequality follows from the concavity of the $\log$ and the second from the definition of $C_{k^{*}}$. The left-hand side is the derivative of $\gamma\rightarrow\sum_{\theta}\lambda_{\theta}^{*}r_{\theta}^{\gamma}(q)$ evaluated at $\gamma=0$. Hence, for all $q\notin C_{k^{*}}$, there exists $\gamma(q)>0$ such that for all $\gamma\leq\gamma(q)$, $\sum_{\theta}\lambda_{\theta}^{*}r_{\theta}^{\gamma}(q)>0$. Now, if we had a uniform $\gamma_{0}>0$ such that $\sum_{\theta}\lambda_{\theta}^{*}r_{\theta}^{\gamma}(q)>0$ for all $q\notin C_{k^{*}}$, for all $\gamma\leq\gamma_{0}$, and Equation (4) was holding in equality, this would prove that no invariant distribution can put strictly positive probability outside the set $C_{k^{*}}$ for all $\gamma$ small enough. For instance, when $C_{k^{*}}$ is the singleton $\{\lambda^{*}\}$, this would say that $\lambda^{*}$ is an invariant distribution of the Markov chain for all $\gamma$ small enough. This is clearly not true in general. What we can show, however, is that the same logic applies to appropriately chosen neighborhoods of $C_{k^{*}}$. Choosing a sequence of such neighborhoods converging to $C_{k^{*}}$ then completes the proof. In simple terms, for small enough $\gamma_{\ell}$, the invariant measure $\pi_{\ell}$ puts most of its mass on $C_{k^{*}}$, but not all of it.
Observation-dependent Weights. Biases such as the self-confirmation bias require the weights of the updating rule (ER) to depend on the realized signal $x_{n}$ at period $n$ and, possibly, the belief at period $n-1$ (see, e.g., Rabin and Schrag (1999)). Now, if each $\gamma_{n}(x,q)$ is an arbitrary function of $(x,q)$, there is little hope in characterizing the belief process. As Epstein et al. (2010) have already shown, there are instances, where the process converges to a Dirac on a wrong state, even when the model is correctly specified. Similarly, as the blue curve in Figure 6 shows, the process may oscillate perpetually. [^12] To draw Figure 6, we have assumed that $X=\{0,1\}$, $\Theta=\{\theta,\theta^{\prime}\}$, $p^{*}(1)=2/3$, $p_{\theta}(1)=9/10=p_{\theta^{\prime}}(0)$, and $\gamma_{n}(x,q)=1$ if $(B(x,q)-q)(1/2-q)>0$, and $1/n$, otherwise, with $q\in[0,1]$ the belief about $\theta$. In words, the agent updates as a Bayesian when the observation contradicts his belief and is conservative, otherwise.
Figure 6: Observation-dependent weights and oscillation
There is, however, a special case we can deal with. Assume that the weights can be written as $\gamma_{n}(x)=\gamma_{n}\times\alpha(x)$, where $\alpha(x)\in(0,1]$ for all $x\in X$. With this formulation, the weights are independent of the current belief. In addition, the relative weights $\gamma_{n}(x)/\gamma_{n}(y)$ are independent of the number $n$ of observations. If the agent reacts more to the observation $x$ than $y$, he continues to do so throughout. With this formulation, we can rewrite the Robbins-Monro algorithm as:
$q_{n+1}-q_{n}=\gamma_{n+1}\left(\widehat{H}(q_{n})+\widehat{U}_{n+1}\right),$
where $\widehat{H}$ is the function from $\Delta(\Theta)$ to $\mathbb{R}^{|\Theta|}$, whose $\theta$-coordinate is
$\widehat{H}_{\theta}(q):=q_{\theta}\left(\sum_{x}p^{*}(x)\alpha(x)\left(\frac{p_{\theta}(x)}{\sum_{\theta^{\prime}}q_{\theta^{\prime}}p_{\theta^{\prime}}(x)}-1\right)\right),$
and $(\widehat{U}_{n})_{n}$ the bounded martingale difference
$\widehat{U}_{n+1}:=\alpha(x_{n+1})B(q_{n},x_{n+1})-\mathbb{E}_{p^{*}}[\alpha(x_{n+1})B(q_{n},x_{n+1})].$
The analysis is then identical to the one in Section 3. In particular, if we denote $p_{\alpha}^{*}(x):=\frac{p^{*}(x)\alpha(x)}{\sum_{x^{\prime}}p^{*}(x^{\prime})\alpha(x^{\prime})}$ the as-if data generating process, it is immediate to see that the predictive process converges almost surely to:
$\arg\max_{p\in{\rm co\,}\mathcal{P}}\sum_{x\in X}p_{\alpha}^{*}(x)\ln p(x).$
In the special case where $p_{\alpha}^{*}\in{\rm co\,}\mathcal{P}$, the predictive process converges to $p_{\alpha}^{*}$, which can be closer to $p^{*}$ than any of the $p_{\theta}$’s are.
Example 4
There are two states $A$ and $B$, two possible observations $a$ and $b$, and $p_{A}(a)>p_{B}(a)$, that is, observing $a$ (resp., $b$) is more likely when the state is $A$ (resp., $B$). Assume that the agent initially believes that $A$ is more likely, i.e., $q_{0}(A)>1/2$ and let $\alpha(a)>\alpha(b)$. This simple setting captures a form of prior dependence, where the agent reacts more to information, that confirms his initial belief. It is an extreme form of prior dependence as the agent continues to react more to observation $a$, even when his current belief $q_{n}$ indicates that $B$ is the most likely state. The as-if data generating process is:
$p^{*}_{\alpha}(a)=\frac{p^{*}(a)\alpha(a)}{p^{*}(a)\alpha(a)+p^{*}(b)\alpha(b)}>p^{*}(a),$
i.e., it is as if the agent learns in an environment where the data generating process is distorted towards $a$, the observation most indicative of $A$. For a concrete example, let $p^{*}(a)=2/3$, $p_{A}(a)=p_{B}(b)=3/4$, $\alpha(a)=3/4$ and $\alpha(b)=1/4$, so that $p^{*}_{\alpha}(a)=6/7$. The agent’s predictive process therefore converges to $p_{A}$, the same limiting predictive process a Bayesian agent would converge to. If, however, the prior dependance is weaker, say $\alpha(a)=2/3$ and $\alpha(b)=1/2$, then the predictive process converges to $(8/11,3/11)$, which is strictly closer to the true data-generating process than $p_{A}$.
So far, we have assumed that the weights are independent of the current belief $q_{n}$. Within the context of Example 4, suppose instead that we write the updating weights as $\gamma_{n+1}\times\alpha(x_{n+1},q_{n})$, with $\alpha(a,q)=\alpha^{+}=\alpha(b,1-q)$ when $q>1/2$, $\alpha(a,1-q)=\alpha^{-}=\alpha(b,q)$ when $q<1/2$, and $1\geq\alpha^{+}>\alpha^{-}\geq 0$. When $q=1/2$, let $\alpha(a,1/2)=\alpha(b,1/2)=\alpha$. Thus, if the agent observes $a$ (resp. $b$) when he currently believes $A$ (resp. $B$) to be the most likely state, he revises his belief towards the Bayesian posterior more than when he observes $b$ (resp., $a$). This updating process continues to be a Robbins-Monro algorithm with the function $\widehat{H}$:
$\widehat{H}(q):=\begin{cases}p^{*}(a)\left[\frac{p_{A}(a)q}{p_{A}(a)q+p_{B}(a)(1-q)}\right]\alpha^{+}+p^{*}(b)\left[\frac{p_{A}(b)q}{p_{A}(b)q+p_{B}(b)(1-q)}\right]\alpha^{-},&q>1/2,\\ p^{*}(a)\left[\frac{p_{A}(a)}{p_{A}(a)+p_{B}(a)}\right]\alpha+p^{*}(b)\left[\frac{p_{A}(b)}{p_{A}(b)+p_{B}(b)}\right]\alpha,&q=1/2,\\ p^{*}(a)\left[\frac{p_{A}(a)q}{p_{A}(a)q+p_{B}(a)(1-q)}\right]\alpha^{-}+p^{*}(b)\left[\frac{p_{A}(b)q}{p_{A}(b)q+p_{B}(b)(1-q)}\right]\alpha^{+},&q<1/2.\\ \end{cases}$
For a concrete example, let $p^{*}(a)=2/3$, $p_{A}(a)=p_{B}(b)=3/4$, $\alpha^{+}=3/4$, $\alpha^{-}=1/4$, and $\alpha=1/2$. The function $\widehat{H}$ is then
$\widehat{H}(q):=\begin{cases}\frac{2}{3}\left[\frac{3q}{1+2q}-q\right]\frac{3}{4}+\frac{1}{3}\left[\frac{q}{3-2q}-q\right]\frac{1}{4}&q>1/2,\\ \frac{1}{24}&q=1/2,\\ \frac{2}{3}\left[\frac{3q}{1+2q}-q\right]\frac{1}{4}+\frac{1}{3}\left[\frac{q}{3-2q}-q\right]\frac{3}{4}&q<1/2.\end{cases}$
We plot the function $\widehat{H}$ in Figure 7. It is immediate to verify that $\widehat{H}$ has three zeroes: $q=0$, $q=3/10$, and $q=1$, the latter two being locally stable.
Figure 7: The function $\widehat{H}$
The local stability of $q=3/10$ and $q=1$ is not surprising. Since $\alpha(x,\cdot)$ is constant in $q$ for all $q<1/2$, the cross-entropy
$\sum_{x\in\{a,b\}}\frac{p^{*}(x)\alpha(x,0)}{\sum_{x^{\prime}\{a,b\}}p^{*}(x^{\prime})\alpha(x^{\prime},0)}\log\left(qp_{A}(x)+(1-q)p_{B}(x)\right)$
is maximized with respect to $q\in[0,1/2)$ in a stable zero of $\widehat{H}$, restricted to that sub-domain. It is immediate to verify that $q=3/10$ is indeed the maximizer. A symmetric argument applies for all $q>1/2$.
To sum up, the main qualitative feature of our analysis continue to hold: non-Bayesian updating can outperform Bayesian updating in misspecified environments.
4.2 Generalized Bayes Rule
A related updating rule is the generalized Bayes rule, where the revised belief $q_{n+1}(\theta)$ is proportional to
$q_{n}(\theta)p_{\theta}(x_{n+1})^{\gamma_{n+1}},$
when the agent observes $x_{n+1}$, $\gamma_{n+1}\in[0,1]$. The generalized Bayes rule is a special case of the Grether rule (GR), captures some form of underreaction to new information and corresponds to Bayes rule when $\gamma_{n+1}=1$. Recent contributions in statistics, e.g., Grünwald and Van Ommen (2017) and Grünwald and Mehta (2020), advocate the use of the generalized Bayes rule, particularly as a method to correct for inconsistencies in Bayesian inference problems.
We now argue that its long-run properties are markedly different from the rule (ER). To ease the discussion, assume that there exists a unique $p_{\theta^{*}}$, which maximizes the cross-entropy $V$, that is, $V(p_{\theta^{*}})>V(p_{\theta})$ for all $\theta\neq\theta^{*}$. When the observations up to period $n+1$ are $(x_{1},\dots,x_{n+1})$, we have that
$\log\left(\frac{q_{n+1}(\theta^{*})}{q_{n+1}(\theta)}\right)$ $=\log\left(\frac{q_{0}(\theta^{*})}{q_{0}(\theta)}\right)+\sum_{i=1}^{n+1}\gamma_{i}\log\left(\frac{p_{\theta^{*}}(x_{i})}{p_{\theta}(x_{i})}\right),$
$\geq\log\left(\frac{q_{0}(\theta^{*})}{q_{0}(\theta)}\right)+$
$n\Big{[}\Bigl{(}\max_{i=1,\dots,n+1}\gamma_{i}\Bigr{)}\frac{1}{n+1}\sum_{i=1}^{n+1}\log p_{\theta^{*}}(x_{i})-\Bigl{(}\min_{i=1,\dots,n+1}\gamma_{i}\Bigr{)}\frac{1}{n+1}\sum_{i=1,\dots,n+1}^{n+1}\log p_{\theta}(x_{i})\Big{]}.$
The strong law of large numbers implies that $\lim_{n\rightarrow+\infty}\frac{1}{n+1}\sum_{i=1}^{n+1}\log p_{\hat{\theta}}(x_{i})=V(p_{\hat{\theta}})$ almost surely for all $\hat{\theta}$, hence the term into bracket converges to a strictly positive number since $V(p_{\theta^{*}})>V(p_{\theta})$. Therefore, $\left(\log\left(\frac{q_{n+1}(\theta^{*})}{q_{n+1}(\theta)}\right)\right)_{n}$ converges to +$\infty$+ ∞ and, consequently, $(q_{n+1}(\theta^{*}))_{n}$ converges to one almost surely. The generalized Bayes rule thus converges to the same degenerated belief as Bayes rule, in sharp contrast with the updating rule (ER).
Thus, despite the similarities of the two rules, their asymptotic properties are fundamentally different. Intuitively, this is because the two rules underreact to new information on two different scales (linear vs. logarithmic), with the updating rule (ER) underreacting more and, consequently, protecting more against model’s misspecification. However, to the best of our knowledge, no axiomatically founded criteria currently exist for systematically ranking updating rules based on their responsiveness to new information. [^13] While we could employ a (pseudo-)metric on the space of distributions over posteriors, such a choice would be largely arbitrary.
4.3 Overreaction
We have assumed that the agent underreacts to news, i.e., $\gamma_{n}\leq 1$. If the agent overreacts to news, the convergence to a limit point or cycle is not guaranteed in general. As Figure 8 suggests, the belief process (the green curve) may be quite chaotic, with the beliefs varying wildly. [^14] To draw Figure 8, we have assumed that $X=\{0,1\}$, $\Theta=\{\theta,\theta^{\prime}\}$, $p^{*}(1)=2/3$, $p_{\theta}(1)=9/10=p_{\theta^{\prime}}(0)$, and $\gamma_{n}=2+1/n$. In instances where $(1-\gamma_{n})q_{n}+\gamma_{n}B(x_{n+1},q_{n})>1$ (resp., $<0$), we let $q_{n+1}=0.99$ (resp., $0.01$). There is therefore little hope for general statements. We conjecture, however, that if $(\gamma_{n})_{n}$ converges to 1, then we recover the convergence to the closest distribution in $\mathcal{P}$ to $p^{*}$.
Figure 8: Overreaction and Oscillation
5 Literature
Two closely related papers are Heidhues et al. (2021) and Frick et al. (2023). Heidhues et al. (2021) consider the following model. In each period $n$, the agent makes a decision $a_{n}$ and observes a signal $q_{n}$, which depends on the decision $a_{n}$ and an (unobservable) external factor $x_{n}$. The authors assume that $x_{n}=\theta+\epsilon_{n}$, where $\theta$ is a fixed parameter and $\epsilon_{n}$ a realization of a normal distribution with mean zero and precision $\rho_{\epsilon}$. (The realizations are independent.) When the agent makes the decision $a_{n}$ and the external factor is $x_{n}$, the signal is $q_{n}=Q(a_{n},x_{n})$, with $Q$ strictly monotonic in $x$ for all $a$. The agent, however, perceives the relationship between decisions, external factors and signals to be $\widehat{Q}$, also strictly monotonic in $x$ for all $a$. The agent is Bayesian and initially believes that the unknown mean $\theta$ is normally distributed with mean zero and precision $\rho_{0}$. Thus, upon observing $q_{n}$ after having chosen $a_{n}$, the agent thinks that the realized external factor is $\widehat{x}_{n}$, the unique solution to $\widehat{Q}(a_{n},x)=q_{n}$, which may differ from the truly realized $x_{n}$. At period $n$, the agent’s expectation $\widehat{\theta}_{n}$ of the mean therefore satisfies:
$\widehat{\theta}_{n}=\widehat{\theta}_{n-1}+\gamma_{n}[\widehat{x}_{n}-\widehat{\theta}_{n-1}],\text{where\;}\gamma_{n}=\frac{\rho_{\epsilon}}{n\rho_{\epsilon}+\rho_{0}}.$
Heidhues et al. (2021) rewrite the evolution of $\widehat{\theta}_{n}$ as a Robbins-Monro algorithm and prove almost sure convergence. (This discussion sidesteps an important issue, namely that, $\widehat{x}_{n}$ depends on the decision made $a_{n}$.) While Heidhues et al. (2021) frame their problem as an active learning problem, they also explain how we can reframe it as a learning problem with non-Bayesian updating. For instance, the model accommodates underreaction for specifications of $Q$ and $\widehat{Q}$ such that $x_{n}>\widehat{x}_{n}>0$ and $x_{n}<\widehat{x}_{n}<0$. The model can also accommodate base-rate neglect and the confirmatory bias. If we interpret any deviation from $x_{n}$ as a departure from Bayesian updating, Bayesian updating then corresponds to $x_{n}=\widehat{x}_{n}$, that is, for all decisions made, the agent’s model is correct (and, thus, learns $\theta$). Unlike our framework, their framework is not flexible enough to cleanly decouple the updating rule from the agent’s misspecification. This is not surprising as non-Bayesian updating in misspecified models was not their focus. We differ in two other important ways. First, we do not restrict attention to a particular family of distributions, such as the normal distributions. Second, we not only prove that the belief process converges almost surely, but also fully characterize the limit set. Heidhues et al. (2021) only prove convergence. In common, we use the ODE method to characterize the limit set of our stochastic approximation algorithm. See Bervoets et al. (2020), Esponda et al. (2021), Gagnon-Bartsch and Bushong (2022) and Danenberg and Fudenberg (2024) for some other recent applications of stochastic approximation in economics and Marcet and Sargent (1989) for an earlier application.
Frick et al. (2023) propose a general model of Bayesian learning, where the true and perceived data generating processes $p_{q}^{*}\in\Delta(X)$ and $p_{\theta,q}\in\Delta(X)$, $\theta\in\Theta$, may depend on the current belief $q\in\Delta(\Theta)$. The model is flexible enough to accommodate some departures from Bayesian updating as well as several models of active and social learning (in the latter case the dependence on the current beliefs is through the actions chosen). In particular, as we have already argued, the updating rule (ER) induces the same posteriors than the Bayesian updating of the “as-if” data generating processes $p^{\gamma}_{\theta,q}\in\Delta(X)$, where $p^{\gamma}_{\theta,q}(x):=(1-\gamma)(\sum_{\theta^{\prime}}q_{\theta^{\prime}}p_{\theta^{\prime}}(x))+\gamma p_{\theta}(x)$ for all $(x,\theta)$, when the current belief is $q$. The results of Frick et al. (2023) are not applicable to our problem, however. First, their model cannot capture weights that depend on the number $n$ of observations, as we do. Second, and most importantly, their results provide conditions for the (in)stability of degenerated beliefs $\delta_{\theta}$, $\theta\in\Theta$. None of their results addresses the stability properties of interior beliefs. For other related papers on learning in misspecified models, see Esponda and Pouzo (2016), Fudenberg et al. (2017), Heidhues et al. (2018), Frick et al. (2020), Fudenberg et al. (2021), Bohren and Hauser (2021), and Lanzani (2022).
Finally, Frick et al. (2024) compare misspecified models in the spirit of Blackwell (1951). More precisely, consider two statistical models $\mathcal{P}^{1}:=(p_{\theta}^{1})_{\theta}$ and $\mathcal{P}^{2}:=(p_{\theta}^{2})_{\theta}$ and assume that the true model is $\mathcal{P}^{*}:=(p^{*}_{\theta})_{\theta}$. The idea is to compare the expected payoff the agent would obtain if his decisions are based on either model. Formally, let $\mathcal{F}$ be a non-empty set of payoff acts $f:\Theta\rightarrow\mathbb{R}$ and $f_{\mathcal{F}}^{i}(x_{1},\dots,x_{n})$ a solution to $\max_{f\in\mathcal{F}}\mathbb{E}[f|(x_{1},\dots,x_{n}),\mathcal{P}^{i}]$, that is, $f_{\mathcal{F}}^{i}(x_{1},\dots,x_{n})$ maximizes the agent’s expected payoff given the observations $(x_{1},\dots,x_{n})$ and the model $\mathcal{P}^{i}$. (The authors assume Bayesian updating, but as already explained, some choices of $\mathcal{P}^{i}$ are equivalent to non-Bayesian updating.) The model $\mathcal{P}^{1}$ outperforms the model $\mathcal{P}^{2}$ if there exists $n^{*}$ such that for all $n\geq n^{*}$, $\mathbb{E}_{\mathcal{P}^{*}}[f_{\mathcal{F}}^{1}(x_{1},\dots,x_{n})]\geq\mathbb{E}_{\mathcal{P}^{*}}[f_{\mathcal{F}}^{2}(x_{1},\dots,x_{n})]$ for all decision problems $\mathcal{F}$. Frick et al. (2024) provide tight conditions for one model to outperform the other. There are no difficulties to adapt their definitions to our framework. It is, however, unclear how to adapt their results, since the updating rule (ER) necessitates a model, which changes with the current belief and the number of observations. Intuitively, though, when $p^{*}\in{\rm co\,}\mathcal{P}$, the updating rule (ER) converges to the true data-generating process, while Bayesian updating does not. For large enough $n$, the updating rule (ER) therefore outperforms Bayesian updating, in the sense of Frick et al. (2024). Even when $p^{*}\notin{\rm co\,}\mathcal{P}$, the updating rule (ER) comes closer to the true data-generating process, hence there are decision problems (e.g., scoring problems), where it again outperforms Bayesian updating.
6 Proofs
6.1 Proof of Theorem 1
The proof uses methods and results from stochastic approximation theory. We refer the reader to Benaïm (1999) for a short exposition and to Borkar (2008) and Kushner and Yin (2003) for book-length treatments.
The roadmap is as follows. We first rewrite the updating rule (ER) as a Robbins-Monro algorithm, with associated ODE $\dot{q}(t)=H(q(t))$. We then analyze the associated ODE. In particular, we show that the cross-entropy map $V$ is Lyapunov function for the associated flow and that the set $E$ of critical points of $H$ can be decomposed into finitely many closed convex (hence, connected) components (the decomposition is unique). By Proposition 6.4 in Benaïm (1999), the (random) limit set of $(q_{n})_{n}$ is included in a closed connected subset of $E$. Finally, we show that the probability to converge to the component that maximizes the cross-entropy is one.
We rewrite the updating rule (ER) as:
$q_{n+1}$ $=q_{n}+\gamma_{n+1}\left[B(q_{n},x_{n+1})-q_{n}\right]$
$=q_{n}+\gamma_{n+1}\left[H(q_{n})+U_{n+1}\right],$
where $(U_{n})_{n}$ is the bounded martingale difference with $U_{n+1}:=B(q_{n},x_{n+1})-\mathbb{E}_{p^{*}}[B(q_{n},\mathbf{x}_{n+1}]$ when the realized signal is $x_{n+1}$, and $H(q_{n}):=\mathbb{E}_{p^{*}}[B(q_{n},\mathbf{x}_{n+1})]-q_{n}$.
The ODE associated with the random process $(q_{n})_{n}$ is
$\dot{q}(t)=H(q(t)).$ (5)
For further reference, note that the component-wise system is given by
$\dot{q}_{\theta}(t)=q_{\theta}(t)(-1+f_{\theta}(q(t))),\;\,\text{with}\;f_{\theta}(q):=\sum_{x}p^{*}(x)\frac{p_{\theta}(x)}{\sum_{\theta^{\prime}}q_{\theta^{\prime}}p_{\theta^{\prime}}(x)},\forall\theta\in\Theta.$ (6)
We first show that the cross-entropy map $V$ grows along the trajectory of the dynamics (5).
Lemma 1
$V$ is a strict Lyapunov function for the flow $\phi_{t}$ associated to (5), meaning that, along any non-stationary solution curve $(q(t))_{t\geq 0}$ we have $\frac{d}{dt}V(q(t))>0$.
Proof of Lemma 1 Note that
$\frac{d}{dt}V(q(t))=\left<\nabla V(q(t)),\dot{q}(t)\right>=G(q(t))$
for all $t\geq 0$, where $G(q):=\sum_{\theta}\frac{\partial V}{\partial q_{\theta}}(q)q_{\theta}(-1+f_{\theta}(q))$.
Since
$\frac{\partial V}{\partial q_{\theta}}(q)=\sum_{x\in X}p^{*}(x)\frac{p_{\theta}(x)}{\sum_{\theta^{\prime}\in\Theta}p_{\theta^{\prime}}(x)q_{\theta^{\prime}}}=f_{\theta}(q),$
we have
$G(q)=\sum_{\theta\in\Theta}q_{\theta}f_{\theta}(q)(-1+f_{\theta}(q))=\sum_{\theta\in\Theta}q_{\theta}\left(-1+f_{\theta}(q)\right)^{2}+\sum_{\theta\in\Theta}q_{\theta}f_{\theta}(q)-\sum_{\theta\in\Theta}q_{\theta}.$
Notice that $\sum_{\theta\in\Theta}q_{\theta}f_{\theta}(q)=1$, and since $\sum_{\theta\in\Theta}q_{\theta}=1$, we have $G(q)\geq 0$, with equality if and only if $f_{\theta}(q)=1$ for all $\theta$ such that $q_{\theta}>0$. Hence, $G(q)\geq 0$ with equality if and only if $q\in E$, which concludes the proof. $\blacksquare$
Lemma 1 implies that the only invariant sets for the flow are connected components of $E$, the set of zeroes of $H$. By Proposition 6.4 in Benaïm (1999), it also implies that $\mathcal{L}(q_{n})$, the (random) limit set of $(q_{n})_{n}$, is included in a closed connected subset of $E$ [^15] The decomposition of a set into its connected components is always unique. We now characterize $E$.
Lemma 2
  • (i)
    The set $E$ is a finite union of disjoint non-empty compact and convex components: there exists $K<+\infty$ such that $E=\bigcup_{k=1}^{K}C_{k}$, where each $C_{k}$ is a non-empty compact convex subset of $\Delta(\Theta)$ and $C_{k}\cap C_{k^{\prime}}=\emptyset$ for all $(k,k^{\prime})$ with $k\neq k^{\prime}$.
  • (ii)
    The cross-entropy $V$ takes the constant value $V(C_{k})$ on each component $C_{k}$, and there exists a unique $k^{*}$ such that $V(C_{k^{*}})>V(q)$ for all $q\notin C_{k^{*}}$.
Proof of Lemma 2 For any non-empty subset $\widehat{\Theta}$ of $\Theta$, we write $F_{\widehat{\Theta}}$ for the subset of $\Delta(\Theta)$:
$F_{\widehat{\Theta}}:=\left\{q\in\Delta(\Theta):\;\,q_{\theta}=0\text{ iff }\;\theta\notin\widehat{\Theta}\right\}.$
In particular, $F_{\Theta}=\mathbf{S}^{*}$ and $F_{\{\theta\}}=\{\delta_{\theta}\}$.
Recall that $q\in E$ if and only if
$q_{\theta}(f_{\theta}(q)-1)=0,$
for all $\theta$. Consider the constrained maximization problem $\max_{q\in F_{\widehat{\Theta}}}V(q)$, that is,
$\max_{q\in\mathbb{R}^{|\Theta|}}V(q),$
subject to $\sum_{\theta\in\widehat{\Theta}}q_{\theta}=1$, $q_{\theta}>0$ for all $\theta\in\widehat{\Theta}$ and $q_{\theta}=0$ for all $\theta\notin\widehat{\Theta}$. If a solution exists, for all $\theta\in\widehat{\Theta}$, the first-order condition with respect $q_{\theta}$ satisfies
$f_{\theta}(q):=\sum_{x}p^{*}(x)\frac{p_{\theta}(x)}{\sum_{\theta^{\prime}}q_{\theta^{\prime}}p_{\theta^{\prime}}(x)}=\lambda,$
where $\lambda$ is the Kuhn-Tucker multiplier associated with $\sum_{\theta\in\widehat{\Theta}}q_{\theta}=1$. (The complementary slackness condition implies that the multipliers associated with $q_{\theta}>0$ for all $\theta\in\widehat{\Theta}$ are all zero.) Multiplying each equation by $q_{\theta}$ and summing over $\theta\in\widehat{\Theta}$, we get
$\sum_{\theta\in\Theta}q_{\theta}f_{\theta}(q)=\lambda\sum_{\theta\in\Theta}q_{\theta}=\lambda.$
Since the left hand-side is equal to $1$, it follows that
$f_{\theta}(q)=1\text{\;\; for all }\theta\in\widehat{\Theta}.$
Therefore, at a maximum, we have
$q_{\theta}(f_{\theta}(q)-1)=0,$
for all $\theta\in\Theta$, i.e., it is a critical point of $H$. Since $V$ is concave, the converse is also true by the Kuhn-Tucker-Karush theorem. It follows that $E$ is the union of the sets $S_{\widehat{\Theta}}:=\arg\max_{q\in F_{\widehat{\Theta}}}V(q)$ over all possible non-empty subsets $\widehat{\Theta}$ of $\Theta$.
Since $V$ is concave, $S_{\widehat{\Theta}}$ is convex for all $\widehat{\Theta}$. We write $V(S_{\widehat{\Theta}})$ for the value taken by $V$ on $S_{\widehat{\Theta}}$. We now construct the decomposition from the sets $S_{\widehat{\Theta}}$.
[…]
We first show that $C_{k}$ is closed: let $(q^{i})_{i}$ be a sequence in $C_{k}$ and suppose that $\lim_{i\rightarrow+\infty}q^{i}=q$, with $q\in F_{\widehat{\Theta}}$. By continuity, $V(q)=V(C_{k})$. Moreover, for $i$ large enough, $q^{i}_{\theta}>0$, for all $\theta\in\widehat{\Theta}$. Assume by contradiction that $q\notin S_{\widehat{\Theta}}$, then there exists $\tilde{q}\in F_{\widehat{\Theta}}$ such that $V(\tilde{q})>V(q)$. There exists a sequence $\tilde{q}^{i}$ such that $\tilde{q}^{i}$ is in the same face as $q^{i}$, and $\lim_{i\rightarrow+\infty}\tilde{q}^{i}=\tilde{q}$. Therefore
$\lim_{i\rightarrow+\infty}V(q^{i})=V(q)<V(\tilde{q})=\lim_{i\rightarrow+\infty}V(\tilde{q}^{i}),$
implying that, for large enough $i$, $V(\tilde{q}^{i})>V(q^{i})$. This contradicts the fact that $q^{i}$ maximizes $V$ on the face it belongs to. Therefore, $q\in S_{\widehat{\Theta}}$ for some $\widehat{\Theta}\in\mathcal{C}_{k}$. Consequently, the collection $\{C_{k}\}$ is composed of closed, connected, pairwise disjoint sets, and
$E=\bigcup_{k=1}^{K}C_{k}.$ (7)
We now argue that $C_{k}$ is convex, for $k=1,...,K$. Define $W:\Delta(X)\rightarrow\mathbb{R}$ as follows:
$p\in\Delta(X)\mapsto W(p):=\sum_{x}p^{*}(x)\log p(x).$
Since $p^{*}$ has full support on $X$, $W$ is strictly concave on $\Delta(X)$. [^17] More generally, if the support of $p^{*}$ is not $X$, the argument remains valid if we restrict $W$ to the support of $p^{*}$. Let $L:\Delta(\Theta)\rightarrow\Delta(X)$ be the linear map $q\mapsto\left(\sum_{\theta}q_{\theta}p_{\theta}(x)\right)_{x\in X}$. By linearity, $L(S_{\widehat{\Theta}})$ is a convex subset of $\Delta(X)$. By strict concavity of $W$, $W$ is constant on $L(S_{\widehat{\Theta}})$ if, and only if, $L(S_{\widehat{\Theta}})$ is a singleton. Since $V$ is constant on $S_{\widehat{\Theta}}$, and $V=W\circ L$, this implies that $L(S_{\widehat{\Theta}})$ is a singleton. Call $p_{\widehat{\Theta}}:=\{L(S_{\widehat{\Theta}})\}\in\Delta(X)$. By continuity of $L$, $p_{\widehat{\Theta}}=p_{\widehat{\Theta}^{\prime}}$ for $(\widehat{\Theta},\widehat{\Theta}^{\prime})\in\mathcal{C}_{k}\times\mathcal{C}_{k}$, so that $L$ is constant on $C_{k}$ (notice that this is stronger than stating that $V$ is constant along $C_{k}$). Let $(q,q^{\prime})\in C_{k}\times C_{k}$, with $q\in S_{\widehat{\Theta}},q^{\prime}\in S_{\widehat{\Theta}^{\prime}}$, and $\lambda\in(0,1)$. Then, if $q^{\prime\prime}:=\lambda q+(1-\lambda)q^{\prime}$, we have $L(q^{\prime\prime})=L\left(\lambda q+(1-\lambda)q^{\prime}\right)=\lambda L(q)+(1-\lambda)L(q^{\prime})=L(C_{k})$. Since $q$ and $q^{\prime}$ are critical points of $H$, we have
$1=f_{\theta}(q)=\sum_{x}p^{*}(x)\frac{p_{\theta}(x)}{L(C_{k})(x)},\;\forall\theta\in\widehat{\Theta},\;\text{ and }\;\,1=f_{\theta}(q)=\sum_{x}p^{*}(x)\frac{p_{\theta}(x)}{L(C_{k})(x)},\;\forall\theta\in\widehat{\Theta}^{\prime}.$
Now, let $\theta$ be such that $q^{\prime\prime}_{\theta}\neq 0$. then either $q_{\theta}\neq 0$ or $q^{\prime}_{\theta}\neq 0$. In other terms, $\theta\in\widehat{\Theta}\cup\widehat{\Theta}^{\prime}$. Since $\lambda\in(0,1)$, $q^{\prime\prime}_{\theta}>0$ if and only if $\theta\in\widehat{\Theta}\cup\widehat{\Theta^{\prime}}$. Moreover,
$f_{\theta}(q^{\prime\prime})=\sum_{x}p^{*}(x)\frac{p_{\theta}(x)}{L(q^{\prime\prime})(x)}=\sum_{x}p^{*}(x)\frac{p_{\theta}(x)}{L(C_{k})(x)}=1,$
for all $\theta\in\widehat{\Theta}\cup\widehat{\Theta^{\prime}}$. Hence, $q^{\prime\prime}$ is a critical point of $H$ in $S_{\widehat{\Theta}\cup\widehat{\Theta}^{\prime}}$ and, consequently, ${\rm co\,}(C_{k})\subseteq E$. It follows that $\cup_{k^{\prime}}C_{k^{\prime}}=E=\cup_{k^{\prime}}{\rm co\,}(C_{k^{\prime}})$. Now, if $C_{k}\neq{\rm co\,}(C_{k})$, then ${\rm co\,}(C_{k})\subseteq\cup_{k^{\prime}\neq k}C_{k^{\prime}}$, a contradiction with the fact that the components are pairwise disjoint. Therefore, $C_{k}={\rm co\,}(C_{k})$, that is, $C_{k}$ is convex.
Finally, since $\arg\max_{q\in\Delta(\Theta)}V(q)$ is a convex set, there exists a unique index $k^{*}\in\{1,...,K\}$ such that $C_{k^{*}}=\arg\max_{q\in\Delta(\Theta)}V(q)$. This concludes the proof. $\blacksquare$
The final step consists in proving that the convergence to any $C_{k}$ with $k\neq k^{*}$ has probability zero. The following observations play a crucial role. Since $C_{k}$ is convex (and not merely connected), the set $\mathcal{C}_{k}$ admits a maximal element, that is, there exists $\widehat{\Theta}(k)\in\mathcal{C}_{k}$ such that $\widehat{\Theta}\subset\widehat{\Theta}(k)$, for all $\widehat{\Theta}\in\mathcal{C}_{k}\setminus\{\widehat{\Theta}(k)\}$. We call $\widehat{\Theta}(k)$ the support of $C_{k}$. By construction, for $k\neq k^{*}$, $\widehat{\Theta}(k)\neq\Theta$. We also have:
Lemma 3
For all $k\neq k^{*}$, the set $\widehat{\Theta}(k^{*})\setminus\widehat{\Theta}(k)$ is non-empty.
Proof of Lemma 3 If it were not the case, we would have $\widehat{\Theta}(k^{*})\subseteq\widehat{\Theta}(k)$. Therefore, $\max_{q\in\Delta(\widehat{\Theta}(k^{*}))}V\leq\max_{q\in\Delta(\widehat{\Theta}(k))}V$ and, thus, $V(C_{k^{*}})\leq V(C_{k})$, a contradiction. $\blacksquare$
We are now ready to state and prove the final step in the proof.
Lemma 4
For all $k\neq k^{*}$, $\mathbb{P}\left(\mathcal{L}(q_{n})\subseteq C_{k}\right)=0.$
Proof of Lemma 4 Recall that $C_{k}=\bigcup_{\widehat{\Theta}\in\mathcal{C}_{k}}S_{\widehat{\Theta}}$. Pick $q^{*}\in F_{\widehat{\Theta}(k^{*})}$, and let
$\Theta_{3}:=\bigcup_{q\in C_{k}}\left\{\theta\in\Theta:\;q^{*}_{\theta}>0,\;q_{\theta}=0\right\}.$
For $\hat{q}\in C_{k}$, define $\Theta_{1}(\hat{q}):=\left\{\theta\in\Theta:\;q^{*}_{\theta}=\hat{q}_{\theta}=0\right\}$, and $\Theta_{2}(\hat{q}):=\{\theta\notin\Theta_{3}:\;\hat{q}_{\theta}>0\}$.
By Lemma 3, there exists $\theta^{*}\in\Theta_{3}$ such that $\hat{q}_{\theta^{*}}=0$ for all $\hat{q}\in C_{k}$. We stress that the sets $\Theta_{1}(\hat{q})$ and $\Theta_{2}(\hat{q})$ depend on the face $\hat{q}$ belongs to.
By strict concavity of $V$ on ${\rm co\,}\{q^{*},\hat{q}\}$, we have $\frac{\partial}{\partial t}V\left(tq^{*}+(1-t)\hat{q}\right)>0$. Hence,
$10<\sum_{\theta\in\Theta}(q^{*}_{\theta}-\hat{q}_{\theta})f_{\theta}(\hat{q})=\sum_{\theta\in\Theta}q^{*}_{\theta}f_{\theta}(\hat{q})-1=\sum_{\theta\in\Theta_{2}(\hat{q})\cup\Theta_{3}}q^{*}_{\theta}f_{\theta}(\hat{q})-1=\sum_{\theta\in\Theta_{2}(\hat{q})}q^{*}_{\theta}+\sum_{\theta\in\Theta_{3}}q^{*}_{\theta}f_{\theta}(\hat{q})-$ 1 .
Thus, $\sum_{\theta\in\Theta_{3}}q^{*}_{\theta}f_{\theta}(\hat{q})>1-\sum_{\theta\in\Theta_{2}(\hat{q})}q^{*}_{\theta}$. Since $\sum_{\theta\in\Theta_{2}(\hat{q})\cup\Theta_{3}}q^{*}_{\theta}=1$,
$\sum_{\theta\in\Theta_{3}}\mu_{\theta}f_{\theta}(\hat{q})>1,\;\,\forall\hat{q}\in C_{k},$
where $\mu_{\theta}:=\frac{q^{*}_{\theta}}{\sum_{\theta\in\Theta_{3}}q_{\theta}^{*}}$, for all $\theta\in\Theta_{3}$.
By compactness of $C_{k}$, there exists $\varepsilon>0$ such that $\sum_{\theta\in\Theta_{3}}\mu_{\theta}f_{\theta}(\hat{q})>1+3\varepsilon$, for all $\hat{q}\in C_{k}$. By continuity of $f_{\theta}$, there exists an open neighborhood $U$ of $C_{k}$ (in $\Delta(\Theta)$) such that
$\inf_{\hat{q}\in U}\sum_{\theta\in\Theta_{3}}\mu_{\theta}f_{\theta}(\hat{q})>1+2\varepsilon.$
Note that, if $q_{\theta}>0$, then $0\leq\frac{B_{\theta}(q,x)}{q_{\theta}}=\frac{p_{\theta}(x)}{\sum_{\theta^{\prime}}q_{\theta^{\prime}}p_{\theta^{\prime}}(x)}\leq\frac{p_{\theta}(x)}{\min_{\theta^{\prime}\in\Theta}p_{\theta^{\prime}}(x)}<+\infty$ since $p_{\theta^{\prime}}(x)>0$ for all $(\theta^{\prime},x)$. [^18] Again, if the support of $p^{*}$ was a strict subset the support of each $p_{\theta^{\prime}}$, the analysis would hold for all $x$ in the support of $p^{*}$. Let $C(\theta):=\max_{x\in X}\frac{p_{\theta}(x)}{\min_{\theta^{\prime}\in\Theta}p_{\theta^{\prime}}(x)}$. If $q_{n,\theta}>0$, it follows that
$1-\gamma_{n+1}+\gamma_{n+1}C(\theta)\geq\frac{q_{n+1,\theta}}{q_{n,\theta}}=1-\gamma_{n+1}+\frac{\gamma_{n+1}}{q_{n,\theta}}B_{\theta}(q_{n},x_{n+1})\geq 1-\gamma_{n+1}.$ (8)
Since
$\frac{q_{n+1,\theta}}{q_{n,\theta}}=1+\gamma_{n+1}\left(-1+f_{\theta}(q_{n})+\frac{U_{n+1,\theta}}{q_{n,\theta}}\right).$ (9)
we can combining (8) and (9) to obtain $C(\theta)-1\geq-1+f_{\theta}(q_{n})+\frac{U_{n+1,\theta}}{q_{n,\theta}}\geq-1$, for all $n$.
From the Stochastic Taylor Theorem (see e.g. Aliprantis and Border (1999), Theorem 17.17 p. 569), for any $n\in\mathbb{N}^{*}$, there exists a random variable $\zeta_{n}$, such that $|\zeta_{n}|\leq\gamma_{n+1}\left(-1+f_{\theta}(q_{n})+\frac{U_{n+1,\theta}}{q_{n,\theta}}\right)$, and
$\log\left(\frac{q_{n+1,\theta}}{q_{n,\theta}}\right)=\gamma_{n+1}\left(-1+f_{\theta}(q_{n})+\frac{U_{n+1,\theta}}{q_{n,\theta}}\right)-\frac{1}{(1+\zeta_{n})^{2}}\gamma_{n+1}^{2}\left(-1+f_{\theta}(q_{n})+\frac{U_{n+1,\theta}}{q_{n,\theta}}\right)^{2}.$
Defining $z_{n}:=\sum_{\theta\in\Theta_{3}}\mu_{\theta}\log q_{n,\theta}$, it follows that
$z_{n+1}-z_{n}$ $=$ $\sum_{\theta\in\Theta_{3}}\mu_{\theta}\log\left(\frac{q_{n+1,\theta}}{q_{n,\theta}}\right)$
$=$ $\sum_{\theta\in\Theta_{3}}\mu_{\theta}\log\left(1+\gamma_{n+1}\left(-1+f_{\theta}(q_{n})+\frac{U_{n+1,\theta}}{q_{n,\theta}}\right)\right)$
$=$ $\gamma_{n+1}\sum_{\theta\in\Theta_{3}}\mu_{\theta}\left(-1+f_{\theta}(q_{n})+\frac{U_{n+1,\theta}}{q_{n,\theta}}\right)-K_{n+1}\gamma_{n+1^{2}},$
where $K_{n+1}:=\frac{1}{(1+\zeta_{n})^{2}}\left(-1+f_{\theta}(q_{n})+\frac{U_{n+1,\theta}}{q_{n,\theta}}\right)^{2}$. Note that for $n$ large enough $|K_{n+1}|\leq K:=2\max\{1,(C(\theta)-1)^{2}\}$. $|\zeta_{n}|\leq 1/2$. Taking the conditional expectation, we get
$\mathbb{E}\left(z_{n+1}\mid\mathcal{F}_{n}\right)-z_{n}=\gamma_{n+1}\sum_{\theta\in\Theta_{3}}\mu_{\theta}\left(-1+f_{\theta}(q_{n})\right)-\gamma_{n+1}^{2}\mathbb{E}\left(K_{n+1}\mid\mathcal{F}_{n}\right)\geq\gamma_{n+1}\varepsilon-K\gamma^{2}_{n+1},$
on the event $\{q_{n}\in U\}$, for large enough $n$. Assume by contradiction that $\mathbb{P}\left(\mathcal{L}(q_{n})\subseteq U\right)>0$. Since
$\left\{\mathcal{L}(q_{n})\subseteq U\right\}\subseteq\bigcup_{N\in\mathbb{N}}\left\{q_{n}\in U,\;\forall n\geq N\right\},$
there exists $\bar{N}\in\mathbb{N}$ such that
$\mathbb{P}\left(q_{n}\in U,\;\forall n\geq\bar{N}\right)>0.$
However, on the event $\left\{q_{n}\in U,\;\forall n\geq\bar{N}\right\}$, we have $\sum_{\theta\in\Theta_{3}}\mu_{\theta}\left(-1+f_{\theta}(q_{n})\right)>\varepsilon$ for all $n\geq\bar{N}$, implying that, for any $n>\bar{N}$,
$\mathbb{E}(z_{n+1})\geq\mathbb{E}(z_{\bar{N}})+\varepsilon\sum_{m=\bar{N}}^{n}\gamma_{n+1}-K\sum_{m=\bar{N}}^{n}\gamma_{n+1}^{2}.$
Thus, on the event $\left\{q_{n}\in U,\;\forall n\geq\bar{N}\right\}$, we have $\lim_{n\rightarrow+\infty}\mathbb{E}(z_{n})=+\infty$, which cannot happen since $z_{n}\leq 0$. Therefore $\mathbb{P}\left(q_{n}\in U,\;\forall n\geq\bar{N}\right)=0$, the desired contradiction. $\blacksquare$
6.2 Proof of Proposition 1
Let $\tilde{q}\in\mathbf{S}^{*}$ be a zero of $H$, that is, for all $\theta\in\Theta$,
$1=f_{\theta}(\tilde{q})=\sum_{x\in X}\frac{p^{*}(x)}{\sum_{\theta^{\prime}}p_{\theta^{\prime}}(x)\tilde{q}_{\theta^{\prime}}}p_{\theta}(x).$
Define $u\in\mathbb{R}^{|X|}$ as $u(x):=\frac{p^{*}(x)}{\sum_{\theta^{\prime}}p_{\theta^{\prime}}(x)\tilde{q}_{\theta^{\prime}}}$. We have
$u,p_{\theta}-p_{\theta^{\prime}}\right>=0,\;\forall\theta,\theta^{\prime}.$
Since the family $\{p_{\theta}\}_{\theta\in\Theta}$ is full, this means that $u$ is orthogonal to all vectors in the tangent space $\{v\in\mathbb{R}^{|X|}:\sum_{x\in X}v_{x}=0\}$, that is, $u$ is proportional to the unit vector $(1,\dots,1)$. Therefore, there exists a real number $\mu$ such that $p^{*}(x)=\mu\sum_{\theta^{\prime}}p_{\theta^{\prime}}(x)\tilde{q}_{\theta^{\prime}}$, for all $x\in X$. Summing over $X$, we obtain that $\mu=1$. Hence $p^{*}=\sum_{\theta\in\Theta}p_{\theta}\tilde{q}_{\theta}$, and $\tilde{q}\in\Lambda$, which completes the proof.
6.3 Proof of Proposition 2
Let $\widehat{\Theta}$ be a non-empty subset of $\Theta$. The argmax of the map $p\in{\rm co\,}\{p_{\theta}\}_{\theta\in\widehat{\Theta}}\mapsto W(p)$ is a singleton, $\{\bar{p}\}$, by convexity of ${\rm co\,}\{p_{\theta}\}_{\theta\in\widehat{\Theta}}$ and strict concavity of $W$. Since the family is free and $|\widehat{\Theta}|\leq|\Theta|\leq|X|$, the map $\mu:q\mapsto\sum_{\theta\in\Theta}q_{\theta}p_{\theta}$ is injective from $F_{\widehat{\Theta}}$ to $\Delta(X)$. Therefore, there exists a unique $\bar{q}\in F_{\widehat{\Theta}}$ such that $\mu(\bar{q})=\bar{p}$.
6.4 Proof of Theorem 3
Recall that Theorem 3 characterizes the long-run properties of the updating process when the updating weights are constant, i.e., $\gamma_{n}=\gamma>0$ for all $n$. The proof borrows and adapts ideas from the recent work of Benaïm and Schreiber (2019).
We consider the family of random processes $\{(q_{n}^{\gamma})_{n}\}$, where $q_{0}^{\gamma}=q_{0}\in\mathbf{S}^{*}$, and $(q_{n}^{\gamma})_{n}$ is recursively given by
$q^{\gamma}_{n+1}=(1-\gamma)q^{\gamma}_{n}+\gamma B\left(q^{\gamma}_{n},x_{n+1}\right).$ (10)
It will be convenient to rewrite the recursive formula (10) component-wise as follows: for $\theta\in\Theta$,
$\frac{q^{\gamma}_{n+1,\theta}}{q^{\gamma}_{n,\theta}}=F_{\theta}^{\gamma}\left(q^{\gamma}_{n},x_{n+1}\right),\text{ where}\;F_{\theta}^{\gamma}(q,x):=(1-\gamma)+\gamma\frac{p_{\theta}(x)}{\sum_{\theta^{\prime}}p_{\theta^{\prime}}(x)q_{\theta^{\prime}}}.$ (11)
Let $P_{\gamma}$ be the transition kernel associated to the Markov chain $(q^{\gamma}(n))_{n}$. Given $h\in\mathcal{C}\left(\mathbf{S}\right)$, the set of continuous maps from $\mathbf{S}$ to $\mathbb{R}$, define $P_{\gamma}h$ as
$q\in\mathbf{S}\mapsto P_{\gamma}h(q):=\mathbb{E}\left[h(q^{\gamma}_{1})\mid q^{\gamma}_{0}=q\right]=\int_{\mathbf{S}}h(\lambda)P_{\gamma}(q,d\lambda).$
Given $\gamma>0$, define the occupation measure up to time $n$ as
$\Pi^{\gamma}_{n}:=\frac{1}{n}\sum_{m=0}^{n-1}\delta_{q^{\gamma}_{m}},\,\text{ where }\;\delta_{q}\,\text{ denotes the Dirac measure at }\,q.$ (12)
We are interested in the asymptotic behavior of $(\Pi^{\gamma}_{n})_{n}$, when $\gamma$ gets small.
Following Benaïm and Schreiber (2019), define
$r^{\gamma}_{\theta}(q):=\sum_{x\in X}p^{*}(x)\log F^{\gamma}_{\theta}(q,x).$
Intuitively, this quantity represents the expected growth of the belief in $\theta$, when current belief is $q$ and observations are distributed according to $p^{*}$.
Lemma 5
For all $\lambda^{*}\in C_{k^{*}}$ and $q\in\mathbf{S}$, the derivative of the mapping $\gamma\mapsto\sum_{\theta\in\Theta}\lambda^{*}_{\theta}r^{\gamma}_{\theta}(q)$ evaluated in zero is non-negative, and equal to zero if, and only if, $q\in C_{k^{*}}$.
Consequently, for any $q\in\mathbf{S}\setminus C_{k^{*}}$, there exists $\gamma_{0}(q)>0$ such that, for all $0<\gamma<\gamma_{0}(q)$ and all $\lambda^{*}\in C_{k^{*}}$, we have $\sum_{\theta\in\Theta}\lambda^{*}_{\theta}r^{\gamma}_{\theta}(q)>0$.
Proof of Lemma 5 Pick $q\in\mathbf{S}$ and $\lambda^{*}\in C_{k^{*}}$. The derivative of $r_{\theta}^{\gamma}(q)$ with respect to $\gamma$, evaluated in $\gamma=0$ is equal to
$\sum_{x\in X}p^{*}(x)\left(\frac{p_{\theta}(x)}{\sum_{\theta^{\prime}\in\Theta}q_{\theta^{\prime}}p_{\theta^{\prime}}(x)}-1\right)$
Hence, the derivative of $\gamma\mapsto\sum_{\theta}\lambda^{*}_{\theta}r_{\theta}^{\gamma}(q)$ evaluated in $\gamma=0$ is equal to
$\sum_{x\in X}p^{*}(x)\left(\frac{\sum_{\theta\in\Theta}\lambda^{*}_{\theta}p_{\theta}(x)}{\sum_{\theta\in\Theta}q_{\theta}p_{\theta}(x)}-1\right)\geq\sum_{x\in X}p^{*}(x)\log\left(\frac{\sum_{\theta\in\Theta}\lambda^{*}_{\theta}p_{\theta}(x)}{\sum_{\theta\in\Theta}q_{\theta}p_{\theta}(x)}\right)\geq 0,$
where the inequality follows from the concavity of the $\log$. Moreover, by definition of $C_{k^{*}}$ (as the component of the critical points, where the cross-entropy is maximized), the right-hand side holds in equality if, and only if, $q\in C_{k^{*}}$. In addition, the quantity $\sum_{\theta\in\Theta}\lambda^{*}_{\theta}r^{\gamma}_{\theta}(q)$ is equal to zero if $\gamma=0$. Therefore, if $q\notin C_{k^{*}}$, there exists $\gamma(q)>0$ such that, for any $0<\gamma<\gamma(q)$,
$\sum_{\theta\in\Theta}\lambda^{*}_{\theta}r_{\theta}^{\gamma}(q)>0.$
This concludes the proof. $\blacksquare$
Remark 1
Note that there does not exist a uniform threshold $\gamma_{0}>0$ on $\mathbf{S}\setminus C_{k^{*}}$: in words, for any $\gamma>0$, there might exist $q\in\mathbf{S}\setminus C_{k^{*}}$ such that $\sum_{\theta\in\Theta}\lambda^{*}_{\theta}r_{\theta}^{\gamma}(q)<0$. However, a direct consequence of this lemma is that, for any open set $V$ such that $C_{k^{*}}\subset V$, there exists $\gamma_{0}>0$ such that, for any $\gamma<\gamma_{0}$
$\inf_{q\in\mathbf{S}\setminus V}\sum_{\theta\in\Theta}\lambda^{*}_{\theta}r_{\theta}^{\gamma}(q)>0.$
An invariant probability measure for $(q^{\gamma}_{n})_{n}$ is a probability measure $\pi$ on $\mathbf{S}$ such that
$\int_{\mathbf{S}}h(q)\pi(dq)=\int_{\mathbf{S}}P_{\gamma}h(q)\pi(dq),\,\text{for all }\;h\in\mathcal{C}\left(\mathbf{S}\right).$ (13)
The set of such invariant probability distributions is denoted $\mathrm{Inv}\left(P_{\gamma}\right)$. It is a compact and convex subset of $\mathcal{P}(\mathbf{S})$. Its extreme points are the so-called ergodic measures of $P_{\gamma}$.
Given $\pi\in\mathcal{P}(\mathbf{S})$ and $\theta\in\Theta$, define the quantity $r^{\gamma}_{\theta}(\pi):=\int_{\mathbf{S}}r^{\gamma}_{\theta}(q)\pi(dq)$. It is the expected growth of the belief in $\theta$ if the current belief is distributed according to $\pi$. Proposition 1 in Benaïm and Schreiber (2019) states the following: if $\pi$ is an ergodic measure (for $P_{\gamma}$) then
$\lim_{n\rightarrow+\infty}\frac{1}{n}\sum_{m=0}^{n-1}\log F_{\theta}^{\gamma}(q^{\gamma}_{m},x_{m+1})=r^{\gamma}_{\theta}(\pi)\;\text{ for }\,\pi-\text{almost every initial condition }\,q_{0}.$
Moreover $r^{\gamma}_{\theta}(\pi)=0$ for any $\theta\in Supp(\pi)$, where
$Supp(\pi):=\left\{\theta\in\Theta:\;\pi(\{q\in\mathbf{S}:q_{\theta}>0\})=1\right\}.$
The intuition behind this is that, under stationary regime, the supported states have a null expected growth.
Lemma 6
Let $\pi^{\gamma}$ be a weak* limit point of $(\Pi^{\gamma}_{n})_{n}$. Then $\pi^{\gamma}\in Inv(P_{\gamma})$, and
$\sum_{\theta}\lambda^{*}_{\theta}r^{\gamma}_{\theta}(\pi^{\gamma})\leq 0,\;\text{ for all }\;\lambda^{*}\in C_{k^{*}}.$
Proof of Lemma 6 Let $\pi^{\gamma}$ be a weak* limit point of the sequence $(\Pi_{n}^{\gamma})_{n}$. By definition, there exists an increasing sequence $n_{k}$ such that, for any $h\in\mathcal{C}\left(\mathbf{S}\right)$,
$\int_{\mathbf{S}}h(q)\pi^{\gamma}(dq)=\lim_{k}\int_{\mathbf{S}}h(q)\Pi^{\gamma}_{n_{k}}(dq)=\lim_{k}\frac{1}{n_{k}}\sum_{m=0}^{n_{k}-1}h(q^{\gamma}_{m})$
and
$\int_{\mathbf{S}}P_{\gamma}h(q)\pi^{\gamma}(dq)=\lim_{k}\frac{1}{n_{k}}\sum_{m=0}^{n_{k}-1}P_{\gamma}h(q^{\gamma}_{m})$
By Lemma 3(i) in Benaïm and Schreiber (2019), for any $g\in\mathcal{C}(\mathbf{S}\times X)$, denoting $\overline{g}(q):=\sum_{x}p^{*}(x)g(q,x)$, we have
$\lim_{n\rightarrow+\infty}\left|\frac{\sum_{m=0}^{n-1}g(q^{\gamma}_{m},x_{m+1})}{n}-\frac{\sum_{m=0}^{n-1}\overline{g}(q^{\gamma}_{m})}{n}\right|=0,\;\text{almost surely}.$ (14)
Let $g(q,x):=h\left(q\circ F_{\gamma}(q,x)\right)$. [^19] We let $q\circ q^{\prime}$ denote the Hadamard product:$(q\circ q^{\prime})_{\theta}:=q_{\theta}q^{\prime}_{\theta}$ Note that $g(q^{\gamma}_{m},x_{m+1})=h\left(q^{\gamma}_{m}\circ F_{\gamma}\left(q^{\gamma}_{m},x_{m+1}\right)\right)=h\left(q^{\gamma}_{m+1}\right)$. Since $P_{\gamma}h(q)=\sum_{x\in X}p^{*}(x)h(q\circ F_{\gamma}(q,x))=\bar{g}(q)$, and using (14), we get
$\lim_{n\rightarrow+\infty}\left|\frac{\sum_{m=0}^{n-1}h(q^{\gamma}_{m+1})}{n}-\frac{\sum_{m=0}^{n-1}P_{\gamma}h(q^{\gamma}_{m})}{n}\right|=0.$
Hence
$\int_{\mathbf{S}}P_{\gamma}h(q)\pi^{\gamma}(dq)=\lim_{k}\frac{1}{n_{k}}\sum_{m=0}^{n_{k}-1}P_{\gamma}h(q^{\gamma}_{m})=\lim_{k}\frac{1}{n_{k}}\sum_{m=0}^{n_{k}-1}h(q^{\gamma}_{m})=\int_{\mathbf{S}}h(q)\pi^{\gamma}(dq),$
and (13) follows.
We now prove the second claim. Using identity (14) with $g(q,x)=\sum_{\theta\in\Theta}\lambda^{*}_{\theta}\log F_{\theta}^{\gamma}(q,x)$ and using the fact that
$\sum_{m=0}^{n-1}\log F_{\theta}^{\gamma}\left(q^{\gamma}_{m},x_{m+1}\right)=\sum_{m=0}^{n-1}\log\frac{q_{m+1,\theta}^{\gamma}}{q_{m,\theta}^{\gamma}}=\log\frac{q_{n,\theta}^{\gamma}}{q_{0,\theta}^{\gamma}},$
we obtain
$\lim_{n\rightarrow+\infty}\frac{\sum_{\theta\in\Theta}\lambda^{*}_{\theta}\left(\log q_{n,\theta}^{\gamma}-\log q_{0,\theta}^{\gamma}\right)-\sum_{m=0}^{n-1}\sum_{\theta\in\Theta}\lambda^{*}_{\theta}\sum_{x\in X}p^{*}(x)\log\left(F_{\theta}^{\gamma}\left(q^{\gamma}_{m},x\right)\right)}{n}=0.$
Note that $\limsup_{n}\frac{\sum_{\theta\in\Theta}\lambda^{*}_{\theta}\log q_{n,\theta}^{\gamma}}{n}\leq 0$ and $\lim_{n\rightarrow+\infty}\frac{\sum_{\theta\in\Theta}\lambda^{*}_{\theta}\log q_{0,\theta}^{\gamma}}{n}=0$. Moreover
$\lim_{k\rightarrow+\infty}\frac{1}{n_{k}}\sum_{m=0}^{n_{k}-1}\sum_{x\in X}p^{*}(x)\log F_{\theta}^{\gamma}\left(q^{\gamma}_{m},x\right)=\int_{\mathbf{S}}\sum_{x\in X}p^{*}(x)\log F_{\theta}^{\gamma}(q,x)\pi(dq)=r^{\gamma}_{\theta}(\pi^{\gamma}).$
As a result, we obtain that $\sum_{\theta\in\Theta}\lambda^{*}_{\theta}r^{\gamma}_{\theta}(\pi^{\gamma})\leq 0$. $\blacksquare$
Remark 2
It is worth noting that, even though $\pi\in\mathcal{P}(\mathbf{S})$ is such that $\pi(\mathbf{S}\setminus C_{k^{*}})>0$, it does not imply that there exists $\gamma_{0}>0$ such that, for all $\gamma<\gamma_{0}$,
$\sum_{\theta\in\Theta}\lambda^{*}_{\theta}r_{\theta}^{\gamma}(\pi)>0.$
Indeed, regardless how small $\gamma$ is, $Supp(\pi)$ might contain points $q$ such that $\sum_{\theta\in\Theta}\lambda^{*}_{\theta}r_{\theta}^{\gamma}(q)<0$.
Importantly, Lemmas 5 and 6 do not allow us to conclude that there exists $\gamma_{0}>0$ such that, for any $0<\gamma<\gamma_{0}$, we have $\pi^{\gamma}(C_{k^{*}})=1$. As a matter of fact, this statement cannot hold, because an element of $\mathrm{Inv}(P_{\gamma})$ cannot be supported by $C_{k^{*}}$ (when $C_{k^{*}}=\{\lambda^{*}\}$, $\delta_{\lambda^{*}}$ is not invariant for $P_{\gamma}$.
Definition 3
The measure $\pi^{*}\in\Delta(\Delta(\Theta))$ is a limiting measure for the updating rule (ER) if there exists a sequence $(\gamma_{\ell},\pi_{\ell})_{\ell\in\mathbb{N}}$ such that:
  1. 1.
    $\gamma_{\ell}\downarrow 0$,
  2. 2.
    $\pi_{\ell}$ is a weak* limit point of $(\Pi_{n}^{\gamma_{\ell}})_{n}$, for all $\ell$,
  3. 3.
    $\lim_{\ell}\pi_{\ell}=\pi^{*}$ for the weak* topology.
In words, the measure $\pi^{*}$ is a limiting measure if we can find a sequence of updating weights $(\gamma_{\ell})_{\ell}$ and a corresponding sequence $(\pi_{\ell})_{\ell}$ of limit points of $(\Pi_{n}^{\gamma_{\ell}})_{n}$, which converges to $\pi^{*}$ for the weak* topology. We then prove the following:
Theorem 3. Let $\pi^{*}$ be a limiting measure for the updating rule (ER). The support of $\pi^{*}$ is included in $C_{k^{*}}$.
Proof of Theorem 3 Let $\pi^{*}$ be a limiting measure and $(\gamma_{\ell},\pi_{\ell})_{\ell\in\mathbb{N}}$ the corresponding sequence of weights and invariant measures mentionned in Definition 3.
Pick any $\lambda^{*}\in C_{k^{*}}$. Given $\gamma>0$, $q\in\mathbf{S}$, define
$R^{*}(\gamma,q):=\sum_{\theta\in\Theta}\lambda^{*}_{\theta}r_{\theta}^{\gamma}(q).$
Note that $R^{*}(0,q)=0$ and, by first point of Lemma 5,
$R^{*}(\gamma,q)=R^{*}(0,q)+\gamma\frac{\partial R^{*}}{\partial\gamma}(0,q)+o(\gamma),\text{ for all }\;q\in\mathbf{S}.$
Let $O^{*}$ be an open subset of $\mathbf{S}$ such that
$\bigcup_{k\neq k^{*}}C_{k}\subseteq O,\;C_{k^{*}}\subseteq O^{*},$
where $O:=\mathbf{S}\setminus\mathrm{Cl}(O^{*})$. (Recall that the components $(C_{k})_{k=1,\dots,k^{*}}$ are compact, connected and disjoint.)
Fix $\varepsilon>0$. Choose $O^{*}$ small enough so that $\left|\sup_{q\in\mathrm{Cl(O^{*})}}\frac{\partial R^{*}}{\partial\gamma}(0,q)\right|\leq\frac{\varepsilon}{2}$. (Recall that the derivative of $R^{*}$ is zero for all $q\in C_{k^{*}}$ and the derivative is continuous. Hence, such a choice is always possible.) This implies that $|R^{*}(\gamma,q)|\leq\epsilon\gamma$ for all $q\in\mathrm{Cl}(O^{*})$ and small enough $\gamma$.
Now define $c:=\frac{1}{2}\inf_{q\in O}\frac{\partial R^{*}}{\partial\gamma}(0,q)>0$. We then have $R^{*}(\gamma,q)>c\gamma$ for all $q\in O$ and small enough $\gamma$.
Finally, for large enough $\ell$,
$R^{*}(\gamma_{\ell},\pi_{\ell})$ $=$ $\int_{q\in\mathrm{Cl}(O^{*})}R^{*}(\gamma_{\ell},q)\pi_{\ell}(dq)+\int_{q\in O}R^{*}(\gamma_{\ell},q)\pi_{\ell}(dq)$
$>$ $-\varepsilon\gamma_{\ell}\pi_{\ell}(\mathrm{Cl}(O^{*}))+c\gamma_{\ell}\pi_{\ell}\left(O\right)$
By Lemma 6, $R^{*}(\gamma_{\ell},\pi_{\ell})\leq 0$, meaning that $\pi_{\ell}(O)<\frac{\varepsilon}{c}\pi_{\ell}(\mathrm{Cl}(O^{*}))\leq\frac{\varepsilon}{c}$ for large enough $\ell$. Since $O$ is open,
$\pi^{*}(O)\leq\liminf_{\ell\rightarrow+\infty}\pi_{\ell}(O)<\frac{\varepsilon}{c}.$
Since this inequality holds for any $\varepsilon>0$, we have $\pi^{*}(O)=0$. In particular, this proves that $\pi^{*}\left(N^{\delta}(C_{k^{*}})\right)=1$, for any $\delta$-neighborhood of $C_{k}^{*}$, which concludes the proof. $\blacksquare$
Remark 3
For any open set $\delta>0$ containing $C_{k^{*}}$, we have $\liminf_{l\rightarrow+\infty}\mu_{l}(N^{\delta}\left(C_{k^{*}})\right)=1$